### v1

# Entropy-Lens: The Information Signature of Transformer Computations

Abstract
Transformer models have revolutionized fields from natural language processing to computer vision, yet their internal computational dynamics remain poorly understood—raising concerns about predictability and robustness. In this work, we introduce Entropy-Lens, a scalable, model-agnostic framework that leverages information theory to interpret frozen, off-the-shelf large-scale transformers. By quantifying the evolution of Shannon entropy within intermediate residual streams, our approach extracts computational signatures that distinguish model families, categorize task-specific prompts, and correlate with output accuracy. We further demonstrate the generality of our method by extending the analysis to vision transformers. Our results suggest that entropy-based metrics can serve as a principled tool for unveiling the inner workings of modern transformer architectures.
footnotetext: Equal contribution.
1 Introduction
Transformer-based architectures (Vaswani et al., 2023) are widely employed as state of the art models in several fields, from machine translation and search engines to DNA analysis and protein research (Devlin, 2018; Khattab and Zaharia, 2020; Ji et al., 2021; Chandra et al., 2023). Their declination in language modeling of large corpora is referred to as large language models (LLMs), and in computer vision as vision transformers (ViTs). Despite their success and ubiquity, transformers’ inner workings remain largely unknown, resulting in unpredictable behaviour (Wei et al., 2022) and reliability concerns (Schroeder and Wood-Doughty, 2025; Huang et al., 2025).
Therefore, considerate research efforts are devoted to transformer-based architectures interpretability, mostly focusing on LLMs (Nanda and Bloom, 2022; Bereska and Gavves, 2024; Elhage et al., 2021, 2022) and ViTs (Chefer et al., 2021). While exciting results have been achieved in this area, they remain limited to toy models and simplified setups, both very different from real use-case conditions. Moreover, these methods often require training a set of probes (nostalgebraist, 2020; Belrose et al., 2023) or full models on ad-hoc tasks (Nanda et al., 2023), making them architecture or even model specific and computationally expensive. These limitations restrict the scope and usability of current methodologies, rendering them unsuitable for off-the-shelf or large-scale transformer-based architectures.
Hence, we develop Entropy-Lens, a scalable framework to address these limitations. Our methodology is architecture agnostic and applicable to frozen off-the-shelf large-scale transformers. In our experiments, we consider both LLMs, including Llama (Touvron et al., 2023), Gemma (Team et al., 2024), GPT (Radford, 2018) up to 9B parameters, and ViTs (Wu et al., 2020) and data-efficient image transformers (DeiTs) (Touvron et al., 2021; Wightman, 2019). In particular, we analyze the evolution of the generated tokens’ Shannon entropy after each intermediate block in the residual stream, as described in Section 4. In Section 5, we show that these quantities constitute an information-theoretic signature of which and how the computation is performed. Finally, our work paves the way for several potential research directions involving information theory for LLM interpretability, as outlined in Section 6.
Our contributions are as follows:
  1. 1.
    We develop a scalable model agnostic methodology grounded in information theory for frozen off-the-shelf large-scale transformer architectures interpretability (Section 4).
  2. 2.
    We analyze the Shannon entropy of LLMs’ generated tokens’ intermediate predictions, showing that these quantities (1) identify the model family that generated them (2) identify the ‘task type’ of the prompt (3) correlate with the correctness of LLM generated answers to multiple choice questions (Section 5).
  3. 3.
    We demonstrate that the same framework is adaptable to domains outside of language modeling. In particular, we apply it to computer vision to show that the Shannon entropy of ViTs’ intermediate predictions identify the model family that generated them (Section 5), similarly to the LLMs case.
2 Related Work
Lenses in LLMs
Mechanistic interpretability (Bereska and Gavves, 2024) aims to provide a precise description and prediction of transformer-based computations. Common tools in the field are lenses, which are a broad class of probes deployed in intermediate steps of the residual stream. For example, logit-lens (nostalgebraist, 2020) uses the model’s decoder function to decode the intermediate activations in the vocabulary space. tuned-lens (Belrose et al., 2023) refines this technique by training a different affine probe at each layer, instead of only using the pretrained model’s decoder function. Building on the Transformer-Lens library (Nanda and Bloom, 2022), we propose Entropy-Lens, which employ logit-lens to study and characterize LLMs’ computations via their decoded version with information theory.
Transformers’ Circuits
Another approach to mechanistic interpretability aims to identify and understand the specific sub-computations, or circuits, in a neural network (Saphra and Wiegreffe, 2024). Olah et al. (2020) pioneered this approach, introducing the concept of circuits and demonstrating their existence in small models through manual analysis. In transformers, circuits are hypothesized to act as agents that read from and write to the residual stream, which acts as a form of memory. This has been demonstrated in a simplified transformer model composed only of attention blocks (without MLPs) (Elhage et al., 2021). However, there is evidence that full transformers exhibit similar behavior. This evidence—linked to the concept of superposition, the idea that models can represent more features than the available dimensions by compressing multiple features into one (Elhage et al., 2022)—is supported by studies on sparse autoencoders, which demonstrate the ability to decompose representations into simpler components (Bricken et al., 2023). Conmy et al. (2023a) developed a toolkit to facilitate mechanistic interpretability, offering techniques like activation patching and weight factorization. Building upon this, Conmy et al. (2023b) explored automated circuit discovery methods, addressing the challenge of scaling analysis to larger models. These works collectively emphasize the importance of understanding the concrete computational steps within LLMs, moving beyond superficial observations to reveal the underlying mechanisms.
Information Theory in Transformers
Information theory has been studied both in connection to the training phase of LLMs and their interpretability. For example, a collapse in attention entropy has been linked to training instabilities (Zhai et al., 2023) and matrix entropy was employed to evaluate “compression” in LLMs (Wei et al., 2024). Additionally, mutual information was used to study the effectiveness of the chain-of-thought mechanism (Ton et al., 2024). Our work, instead, shifts the focus on the vocabulary’s natural domain. Through Entropy-Lens, we use information theory to study the evolution of entropy of the intermediate layers’ decoded logits.
3 Background
3.1 Information Theory
The main information-theoretic quantity used in our study is entropy. Given a discrete [^1] Although entropy can be naturally extended to the continuous case with probability density functions, we restrict ourselves to the discrete case as it is the most relevant to our study. random variable $X$ with outcomes $x_{i}$ and probability mass function $p$, the Shannon entropy $H$ of $X$ is defined as
$H(X)=-\sum_{i}p(x_{i})\log p(x_{i})=\mathbb{E}[-\log p(X)]$ (1)
Shannon proved that this function is the only one—up to a scalar multiplication—that satisfies intuitive properties for measuring ‘disorder’. These include being maximal for a uniform distribution, minimal for the limit of a Kronecker delta function, and ensuring that $H(A,B)\leq H(A)+H(B)$ for every possible event $A$ and $B$. The same function already existed in continuous form in physics, where it linked the probabilistic formalism of statistical mechanics with the more phenomenological framework of thermodynamics, where the term ‘entropy’ was originally coined.
Next, we study the entropy of vocabulary predictions—a quantity that is maximal when the prediction assigns equal probability to all tokens, minimal when it assigns zero probability to all but one token, and takes intermediate values when probability is distributed across multiple tokens, consistent with the previously mentioned properties.
3.2 The Transformer
Architecture
Figure 1: Structure of a generic Transformer block.
The transformer (Vaswani et al., 2023) is a deep learning architecture widely applied in language modelling with LLMs (Brown et al., 2020) and computer vision (Dosovitskiy et al., 2021). Transformer computations happen through transformer blocks and residual connections, as exemplified in Figure 2. While various design choices are possible, blocks are usually a composition of layer normalization (Zhang and Sennrich, 2019), attention, and multi layer perceptrons (MLPs), as shown in Fig. 1. Residual connections, instead, sum the output of the layer $i-1$ to the output of the layer $i$.
Inside a single transformer block, the information flows both horizontally and vertically. The former, enabled by the attention mechanism, allows the token representations to interact with each other. In a language modelling task, for example, this is useful to identify which parts of the input sequence—the sentence prompt—should influence the next token prediction and quantify by how much. The latter vertical information flow allows the representation to evolve and encode different meanings or concepts. Usually, the dimension of the latent space is the same for each block in the transformer. The embedding spaces where these computations take place are generally called the residual stream.
Computation schema
LLMs are trained to simply predict the next token in a sentence. That is, given a sentence prompt $S$ with tokens $t_{1},\dots,t_{N}$, the transformer encodes each token with a linear encoder $E$. Throughout the residual stream, the representation $x_{N}$ of the token $t_{N}$ evolves into the representation of the token $t_{N+1}$, which is then decoded back into token space via a linear decoder $D$, sometimes set to $E^{\top}$, tying the embedding weights and the decoder. Finally, the logits—the output of $D$—are normalized with $\operatorname{softmax}$ to represent a probability distribution over the vocabulary. We summarize this operation with the function $W:=\operatorname{softmax}\circ\,D$.
In formal terms, information processing can be expressed using the encoder, decoder, Transformer block $f$, and residual connection:
$x_{j}^{0}$ $=E(t_{j})$ (2)
$x_{j}^{i}$ $=f^{i}(x_{j}^{i-1})+x_{j}^{i-1}$ (3)
$y_{j}^{i}$ $=W(x_{j}^{i})$ (4)
where $j\in\{1,\dots,N\}$ ranges over the number of tokens in the prompt and $i\in\{0,\dots,L\}$ ranges over the number of layers. Hence, $x_{j}^{i}$ represents the activations of token $t_{j}$ after layer $i$.
3.2.1 Instruct Models
Training a Large Language Model (LLM) requires vast amounts of data and is generally divided into multiple phases.
Pretraining: The model is exposed to large datasets through self-supervised tasks, such as next-token prediction or similar variants. This phase helps the model learn a broad range of general knowledge.
Fine-tuning: This phase teaches the model to generate more useful and coherent responses. Two main strategies are used: Chat: The model is trained on structured conversations between a user and the model, with clearly defined roles. Instruct: The model learns from simple commands, without a predefined dialogue structure.
RLHF (optional): Some models undergo Reinforcement Learning from Human Feedback (RLHF) to further refine their responses based on human preferences.
For our experiments, we used off-the-shelf models without RLHF to analyze information processing in a less biased transformer version. We also focused on Instruct models instead of Chat models for two reasons: 1. the Instruct strategy aligns better with our experimental setup 2. Instruct models are more flexible and often preferred for practical applications.
4 Method
The aim of our framework is to find and characterize the information-theoretic signature of transformer computations. Entropy-Lens’s pipeline comprises three steps and is described in Figure 2.
Notation
We denote the input sentence comprising tokens $t_{1},\dots,t_{N}$ by $S=(t_{i})_{i=1}^{N}$. Then, $x_{j}^{i}$ denotes the activations of the token $t_{j}$ after block $i$ for $j\in\{1,\dots,N\}$ and $i\in\{1,\dots,L\}$. Since our analysis focuses on the logits extracted from the intermediate layers of the transformer, it will be useful to distinguish between normalized and unnormalized logits. We define $W:=\operatorname{softmax}\circ\,D$ and $y_{j}^{i}:=W(x_{j}^{i})$ the normalized logits of the token $t_{j}$’s activations after layer $i$.
The core of our methodology is to analyze the entropy of the generated tokens’ intermediate representations $y_{j}^{i}$. These vectors are probability distributions, as they are the output of a $\operatorname{softmax}$. To obtain a single quantity that summarizes the information they contain, we compute their entropy $H(y_{j}^{i})$. For one generated token, we can consider the entropy of all of its intermediate predictions $H(y_{j}^{i})$ for $i\in\{1,\dots,L\}$. This leads us to the definition of entropy profile:
Defintion 1 (Entropy profile)
Let $h_{j}^{i}=H(y_{j}^{i})$ be the entropy of the intermediate representation of token $t_{j}$ after block $i$ and residual connection. The entropy profile of the next generated token is defined as
$h_{N}=\bigoplus_{i}h_{N}^{i}$ (5)
where $\bigoplus$ denotes any aggregation function.
In our experiments, we set $\bigoplus$ to be concatenation, so that $h_{N}=(h_{N}^{1},\dots,h_{N}^{L})^{\top}$, but other choices are possible. The extraction of entropy profiles is the step 1 of our pipeline.
Then, we fix the number of tokens that the LLM is required to generate, $T$ and repeat the same procedure for each of them, leading us to the next definition:
Defintion 2 (Aggregated entropy profile)
Let $h_{N+t}$ be the entropy profiles according to Definition 1 for $t\in\{0,\dots,T-1\}$, i.e.  the entropy profile of each token generated sequentially by a transformer. The aggregated entropy profile of the next $T$ generated tokens is defined as
$h_{[N:T]}=\bigotimes_{t=0}^{T-1}h_{N+t}$ (6)
where $\bigotimes$ denotes any aggregation function.
Note that $\bigotimes$ in Definition 2 need not be the same as $\bigoplus$ defined in Definition 1. In our experiments, we set both of them to be concatenation, so that $h_{[N:T]}$ is the matrix with $h_{N+t}$ as columns, that is $(h_{[N:T]})_{t}^{i}=h_{N+t}^{i}$ for $i\in\{1,\dots,L\}$ and $t\in\{0,\dots,T-1\}$. The aggregation of entropy profiles is the step 2 of our framework.
The last step of our framework is classification, where we feed the aggregated entropy profile to a classifier $\mathcal{C}$ to determine whether it contains sufficient information to identify a particular ‘entity’.
In our experiments, we examine whether aggregated entropy profiles identify model family (Section 5.1.1), task type (Section 5.1.2), and correct and wrong answers to multiple choice questions (Section 5.1.3) in LLMs. Additionally, we apply the same pipeline to ViTs showing the flexibility of the proposed methodology (Section 5.2). In our experiments, we take $\mathcal{C}$ to be a k-NN classifier. Classification of the aggregated entropy profiles is the step 3 our framework.
Figure 2: Entropy-Lens’s pipeline. (Left) A diagram representing a transformer architecture: hidden representations are converted into intermediate predictions with $W$ before calculating their entropy $H$. (Right) A diagram representing our framework: step 1: entropy profile extraction, step 2: entropy profile aggregation and step 3: classification.
5 Experiments
The goal of our experiments is to demonstrate the effectiveness of entropy profiles in distinguishing computational signatures across various models and scenarios. In the first part, we explore how entropy profiles can differentiate between LLM families and characterize distinct tasks. In the second part, we apply our framework to the Vision Transformer architecture to assess whether insights from the LLM experiments can be similarly recovered.
5.1 Entropy-Lens for Large Language Models
The experiments on LLMs focus on three key aspects. First, we demonstrate that entropy profiles are indicative of model families, with distinctions becoming more pronounced as model size increases. Second, we investigate whether the entropy profile of a model alone can be used to classify the task it is performing. Finally, we extend this analysis to assess whether entropy profiles can distinguish between correct and incorrect task execution.
5.1.1 Entropy profiles identify model families
We assess whether aggregated entropy profiles can distinguish different model families by visualizing and analyzing those of 9 models from 3 different families (GPT, Gemma and LLama) with parameter counts ranging from 100M to 9B.
We compute the mean of the entropy profiles obtained from 64 generated tokens, obtained with the prompt ‘The concept of entropy, a brief essay:’, as shown in Figure 3. We observe that the profiles relate uniquely to the model family, rather than a particular model, independently of its size.
The GPT model class starts with high vocabulary entropy in the early layers, indicating a wide range of possible response tokens. Then, entropy gradually decreases—more smoothly than in other classes—leading to a low-entropy state, where the model narrows down to a small set of possible response tokens.
The Gemma model class, on the other hand, starts with low entropy in the very first layers, then rises to higher entropy in the intermediate layers, and finally decreases to low entropy again just before the last layers, where the model is required to produce an output token.
The Llama model class follows a similar pattern, but with a steeper rise, resulting in a higher entropy value maintained over a larger range of intermediate layers.
We observe that the equivalence between models of the same family but different sizes holds when looking at the entropy trend not as a function of the absolute layer index, but rather as the relative layer position within the model.
We conjecture that high entropy phases, whether in the early or intermediate layers, allow the model to explore more possibilities in its response, similarly to how temperature helps avoid getting stuck in local minima in optimization. Then, at the moment of selection, the distribution is ‘cooled down’, forcing the output to be limited to a few possible tokens.
Figure 3: Average entropy profiles over 64 generated tokens per model. The $x$ axis is normalized for an easier comparison when models have a different number of layers.
5.1.2 Entropy profiles identify task types
We verify whether the entropy profiles can identify task types examining generative (continue a text), syntactic (count the number of words in a text), and semantic (extract the subject or moral of a text) tasks.
We do this with the TinyStories dataset (Eldan and Li, 2023). For evaluation robustness, we construct for each task type three prompt templates using a combination of task-specific task prompts, reported in Appendix A Table 3, and a story from TinyStories. These templates are:
  • Base, of the form task prompt + story
  • Reversed, of the form story + task prompt
  • Scrambled, of the form task prompt + scrambled story or scrambled story + task prompt, at random. A scrambled story is a ‘story’ obtained by randomly shuffling the words in a given story from TinyStories.
Note that, for a robust evaluation, we also use 2 possible task prompt variations, as per Table 3.
We generate 800 prompts per task type, $1/3$ of them with the base template, $1/3$ with the reversed template, and $1/3$ with the scrambled templates, for a total of 2400 prompts. We then apply our pipeline from Section 4 to classify the aggregated entropy profiles of these prompts against their task type using a k-NN classifier. The model was evaluated in a 10-fold cross-validation using the ROC-AUC score (one-vs-rest). Table 1 shows the results obtained for 6 models with parameter counts ranging from 1B to 9B. Figure 4 shows the average entropy profiles per task type.
We observe that all k-NN classifiers (i.e. one for each LLM) achieve high accuracy in distinguishing entropy profiles, with a trend toward improved performance for larger models.
Figure 4: Average entropy profiles with shaded standard deviation for different task types: generative (continuation_prompt 1 and 2), syntactic (counting_prompt 1 and 2), and semantic (semantic_prompt 1 and 2). These tasks are induced with the prompts described in Appendix A. Left: Llama-3.2-Insruct. Right: Gemma-2-Instruct
Table 1: ROC-AUC (one-vs-rest) results of different models on the TinyStories task classification. The standard deviation is calculated on the 10-fold cross-validation splits.
Model Model Size k-NN ROC-AUC
Gemma-2-it 2.1B 97.66 ± 0.47
Gemma-2-it 8.9B 98.38 ± 0.50
Llama-3.2-Instruct 1 B 94.94 ± 0.79
Llama-3.2-Instruct 3 B 94.77 ± 0.93
Llama-3-Instruct 8 B 96.10 ± 0.67
Phi-3 3.6B 97.07 ± 0.87
5.1.3 Entropy profiles identify correct task execution
We test whether entropy profiles can identify correct and wrong answers generated by LLMs using the Massive Multitask Language Understanding (MMLU) dataset Hendrycks et al. (2021).
MMLU consists of multiple-choice questions across 57 subjects, ranging from history and physics to law, mathematics, and medicine. The difficulty levels span from elementary to professional, making it a benchmark for evaluating language models on specialized domains. Each dataset entry contains: a question string, four answer choices and a label indicating the correct answer.
We evaluated two models, a Llama-3.2 with 3B parameters Instruct and a Gemma-2 with 2B parameters, by presenting the multiple-choice questions in three different formats (as per Table 4 in Appendix A):
  • Base: A minimal version containing the topic, the question, and multiple-choice answers.
  • Instruct: A version with a brief explanation that it’s a multiple-choice test where only one option should be selected.
  • Humble: A version that also instructs the model to pick a completely random option if it doesn’t know the answer.
Then, we applied our pipeline to extract and aggregate and the responses’ entropy profiles and classify them against the correctness of the corresponding LLM-generated answer. We train a k-NN classifier for each LLM and validate it using 10-fold cross-validation. We also conducted a t-test to compare our classifier to a dummy model. This dummy model generates predictions randomly, sampled from a distribution that reflects the proportion of correct and incorrect answers produced by the LLM, ensuring robustness against class imbalance. The results reject the null hypothesis with the k-NN achieving an AUC-ROC between 67.23 and 73.61, depending on prompt type and model (Table 2).
We observe that the instruct and humble prompts improve Llama’s average accuracy, while for Gemma, this is only true for the instruct prompt. Additionally, in Llama, the model’s higher accuracy seems to be partially linked to greater difficulty in distinguishing correct from incorrect entropy profiles, though more rigorous analysis is needed to confirm this. In Gemma, however, this claim is harder to support.
Figure 5: Entropy profiles for ViT model families.
5.2 Entropy-Lens for Vision Transformers
Finally, to demonstrate the versatility and robustness of our approach beyond language modeling, we analyze the entropy profiles of ViTs and DeiTs.
Using 20 classes from ImageNet-1K (Russakovsky et al., 2015), with 20 images per class, and without any modifications to our framework, we generate the entropy profiles shown in Figure 5. We observe that all profiles start with high entropy values, which then decrease, mostly in the final layers. This behavior is qualitatively similar to that of GPTs or larger LLaMa models (Section 5.1.1), suggesting a universal pattern across domains as different as image processing and natural language processing.
Focusing on computer vision models, we note that while ViT and DeiT families exhibit qualitatively similar trends, they differ quantitatively—ViTs start with higher entropy values, making them easily distinguishable from DeiTs.
Notably, the only profile that stands out is that of ViT Large (with $\sim 300$M parameters), compared to the other models analyzed in this section, which have $\leq 86$M parameters.
For ViT Large, entropy decreases more smoothly, appearing not only as a better approximation of the sharp drop seen in smaller models but possibly following a different behavior entirely, with the entropy decline starting earlier.
We hypothesize a phase transition in entropy behavior as model size increases, occurring somewhere between $87$M and $307$M parameters.
Table 2: Performance comparison of different models and prompt types on MMLU accuracy and k-NN AUC-ROC score.
Model Prompt Type LLM Accuracy k-NN AUC-ROC
Llama-3.2-3B-Instruct Base 50.89 73.61 ± 1.52
Humble 58.51 69.90 ± 1.06
Instruct 60.62 67.23 ± 1.62
Gemma-2-2B-it Base 56.10 71.88 ± 1.63
Humble 54.71 72.78 ± 1.15
Instruct 56.38 68.36 ± 1.23
6 Conclusions
In this work, we prototyped a novel model-agnostic interpretability framework for large-scale transformer-based architectures grounded in information-theory. In Section 5.1.1, we showed that the entropy profiles of LLM intermediate predictions identify the LLM’s model family. In Section 5.2 we conduct similar experiments on vision transformers, demonstrating the wide applicability of our framework. Additionally, in Section 5.1.2 we showed that the same entropy profiles can be used to identify the ‘task type’ in LLMs, and in Section 5.1.3 we used them to distinguish between correct and wrong LLM generated answers. Importantly, all our experiments were conducted on frozen off-the-shelf large-scale transformers.
Appendix A Evaluation Details
Table 3: Prompt templates used for TinyStories tasks.
Task Type Task prompt
Generative How can the story be continued?
Which could be a continuation of the story?
Syntactic How many words are in the story?
Count the number of words in the story.
Semantic What is the main idea of the story?
Who is the subject of the story?
Table 4: Prompt templates used for the MMLU dataset.
Prompt Type Prompt
Base
Subject: {subject}
Question: {question}
Choices:
A. {option_1}
B. {option_2}
C. {option_3}
D. {option_4}
Answer:
Instruct
The following is a multiple-choice question about {subject}. Reply only with the correct option.
Question: {question}
Choices:
A. {option_1}
B. {option_2}
C. {option_3}
D. {option_4}
Answer:
Humble
The following is a multiple-choice question about {subject}. Reply only with the correct option.
If you are unsure about the answer, reply with a completely random option.
Question: {question}
Choices:
A. {option_1}
B. {option_2}
C. {option_3}
D. {option_4}
Answer:

### v2

@@
 Abstract
-Transformer models have revolutionized fields from natural language processing to computer vision, yet their internal computational dynamics remain poorly understood—raising concerns about predictability and robustness.
-In this work, we introduce Entropy-Lens, a scalable, model-agnostic framework that leverages information theory to interpret frozen, off-the-shelf large-scale transformers.
-By quantifying the evolution of Shannon entropy within intermediate residual streams, our approach extracts computational signatures that distinguish model families, categorize task-specific prompts, and correlate with output accuracy.
-We further demonstrate the generality of our method by extending the analysis to vision transformers.
-Our results suggest that entropy-based metrics can serve as a principled tool for unveiling the inner workings of modern transformer architectures.
-footnotetext: Equal contribution.
+Transformer models map input token sequences to output token distributions, layer by layer.
+While most interpretability work focuses on internal latent representations, we study the evolution of these token-level distributions directly in vocabulary space.
+However, such distributions are high-dimensional and defined on an unordered support, making common descriptors like moments or cumulants ill-suited.
+We address this by computing the Shannon entropy of each intermediate predicted distribution, yielding one interpretable scalar per layer.
+The resulting sequence—the entropy profile—serves as a compact, information-theoretic signature of the model’s computation.
+We introduce Entropy-Lens, a model-agnostic framework that extracts entropy profiles from frozen, off-the-shelf transformers.
+We show that these profiles (i) reveal family-specific computation patterns invariant under depth rescaling, (ii) are predictive of prompt type and task format, and (iii) correlate with output correctness.
+We further show that Rényi entropies yield similar results within a broad range of $\alpha$ values, justifying the use of Shannon entropy as a stable and principled summary.
+Our results hold across different transformers, without requiring gradients, fine-tuning, or access to model internals.
+00footnotetext: Preprint - This manuscript is currently under review.11footnotetext: Code available at: github.com/christopher-irw/Entropy-Lens
 1 Introduction
-Transformer-based architectures (Vaswani et al., 2023) are widely employed as state of the art models in several fields, from machine translation and search engines to DNA analysis and protein research (Devlin, 2018; Khattab and Zaharia, 2020; Ji et al., 2021; Chandra et al., 2023).
-Their declination in language modeling of large corpora is referred to as large language models (LLMs), and in computer vision as vision transformers (ViTs).
-Despite their success and ubiquity, transformers’ inner workings remain largely unknown, resulting in unpredictable behaviour (Wei et al., 2022) and reliability concerns (Schroeder and Wood-Doughty, 2025; Huang et al., 2025).
-Therefore, considerate research efforts are devoted to transformer-based architectures interpretability, mostly focusing on LLMs (Nanda and Bloom, 2022; Bereska and Gavves, 2024; Elhage et al., 2021, 2022) and ViTs (Chefer et al., 2021).
-While exciting results have been achieved in this area, they remain limited to toy models and simplified setups, both very different from real use-case conditions.
-Moreover, these methods often require training a set of probes (nostalgebraist, 2020; Belrose et al., 2023) or full models on ad-hoc tasks (Nanda et al., 2023), making them architecture or even model specific and computationally expensive.
-These limitations restrict the scope and usability of current methodologies, rendering them unsuitable for off-the-shelf or large-scale transformer-based architectures.
-Hence, we develop Entropy-Lens, a scalable framework to address these limitations.
-Our methodology is architecture agnostic and applicable to frozen off-the-shelf large-scale transformers.
-In our experiments, we consider both LLMs, including Llama (Touvron et al., 2023), Gemma (Team et al., 2024), GPT (Radford, 2018) up to 9B parameters, and ViTs (Wu et al., 2020) and data-efficient image transformers (DeiTs) (Touvron et al., 2021; Wightman, 2019).
-In particular, we analyze the evolution of the generated tokens’ Shannon entropy after each intermediate block in the residual stream, as described in Section 4.
-In Section 5, we show that these quantities constitute an information-theoretic signature of which and how the computation is performed.
-Finally, our work paves the way for several potential research directions involving information theory for LLM interpretability, as outlined in Section 6.
-Our contributions are as follows:
-  1. 1.
-    We develop a scalable model agnostic methodology grounded in information theory for frozen off-the-shelf large-scale transformer architectures interpretability (Section 4).
-  2. 2.
-    We analyze the Shannon entropy of LLMs’ generated tokens’ intermediate predictions, showing that these quantities (1) identify the model family that generated them (2) identify the ‘task type’ of the prompt (3) correlate with the correctness of LLM generated answers to multiple choice questions (Section 5).
-  3. 3.
-    We demonstrate that the same framework is adaptable to domains outside of language modeling.
-In particular, we apply it to computer vision to show that the Shannon entropy of ViTs’ intermediate predictions identify the model family that generated them (Section 5), similarly to the LLMs case.
+Transformer-based architectures (Vaswani et al.
+2023) operate by iteratively mapping token inputs to distributions over next-token outputs, layer by layer (Shan et al.
+2024).
+This view emphasizes their inherently probabilistic nature: each layer produces a distribution over the vocabulary, encoding the model’s beliefs at that stage.
+Despite their success across domains—from language to biology to vision (Devlin 2018; Ji et al.
+2021; Wu et al.
+2020)—the internal evolution of these distributions remains poorly understood, resulting in unpredictable behaviour (Wei et al.
+2022) and reliability concerns (Schroeder and Wood-Doughty 2025; Huang et al.
+2025).
+Most interpretability research either focuses on toy models (Elhage et al.
+2021, 2022) or requires training a set of probes (nostalgebraist 2020; Belrose et al.
+2023) or full models on ad-hoc tasks (Nanda et al.
+2023), limiting scalability and generality.
+Moreover, existing methods primarily study latent representations—hidden states, attention maps, or MLP activations (Nanda and Bloom 2022; Bereska and Gavves 2024; Chefer, Gur, and Wolf 2021)—rather than the output distributions themselves.
+Figure 1: TSNE of the aggregated entropy profiles of different model families using unprompted models. The size of each point is scaled based on the number of parameters of the model.
+We propose to shift perspective: instead of analyzing the internal geometry of latent activations, we focus on the evolution of the model’s token-level predictions across layers.
+These predictions, obtained by decoding intermediate representations via the model’s output head, form probability distributions over the vocabulary.
+Studying their evolution offers a direct, vocabulary-grounded view of the computation.
+However, these distributions pose two challenges: they are high-dimensional (one value per token) and defined over an unordered support (the vocabulary).
+This makes classical descriptors like variance or higher-order cumulants unstable or ill-defined.
+We address both issues by computing the Shannon entropy of each distribution—a scalar, interpretable quantity invariant to token permutation, and reflective of the model’s uncertainty at each layer.
+We introduce Entropy-Lens, a simple, scalable, and model-agnostic framework for studying transformer computations through the evolution of entropy across layers.
+In our experiments, we consider several LLMs, including Llama (Touvron et al.
+2023), Gemma (Team et al.
+2024), and GPT, (Radford 2018) up to 9B parameters.
+For each generated token, we compute the entropy of its intermediate predictions, as described in Section 4, yelding one interpretable value per layer: this constitutes the entropy profile.
+We then aggregate these profiles across tokens and study whether they capture structural information about the internal LLM’s computation (Figure 1).
+In Section 5, we show that: (i) entropy profiles reveal family-specific signatures of transformer architectures, invariant under depth rescaling; (ii) they are predictive of prompt semantics and format; (iii) they correlate with output correctness; (iv) and results are stable across Rényi entropy variants, justifying our use of Shannon entropy as a principled default.
+Entropy-Lens requires no gradients, no fine-tuning, and no access to internal weights.
+It applies to frozen, off-the-shelf transformers of arbitrary size.
+While our experiments focus on language models, the methodology is general and opens new directions for vocabulary-grounded interpretability.
 2 Related Work
 Lenses in LLMs
-Mechanistic interpretability (Bereska and Gavves, 2024) aims to provide a precise description and prediction of transformer-based computations.
+Mechanistic interpretability (Bereska and Gavves 2024) aims to provide a precise description and prediction of transformer-based computations.
 Common tools in the field are lenses, which are a broad class of probes deployed in intermediate steps of the residual stream.
-For example, logit-lens (nostalgebraist, 2020) uses the model’s decoder function to decode the intermediate activations in the vocabulary space.
-tuned-lens (Belrose et al., 2023) refines this technique by training a different affine probe at each layer, instead of only using the pretrained model’s decoder function.
-Building on the Transformer-Lens library (Nanda and Bloom, 2022), we propose Entropy-Lens, which employ logit-lens to study and characterize LLMs’ computations via their decoded version with information theory.
-Transformers’ Circuits
-Another approach to mechanistic interpretability aims to identify and understand the specific sub-computations, or circuits, in a neural network (Saphra and Wiegreffe, 2024).
-Olah et al.
-(2020) pioneered this approach, introducing the concept of circuits and demonstrating their existence in small models through manual analysis.
-In transformers, circuits are hypothesized to act as agents that read from and write to the residual stream, which acts as a form of memory.
-This has been demonstrated in a simplified transformer model composed only of attention blocks (without MLPs) (Elhage et al., 2021).
-However, there is evidence that full transformers exhibit similar behavior.
-This evidence—linked to the concept of superposition, the idea that models can represent more features than the available dimensions by compressing multiple features into one (Elhage et al., 2022)—is supported by studies on sparse autoencoders, which demonstrate the ability to decompose representations into simpler components (Bricken et al., 2023).
-Conmy et al.
-(2023a) developed a toolkit to facilitate mechanistic interpretability, offering techniques like activation patching and weight factorization.
-Building upon this, Conmy et al.
-(2023b) explored automated circuit discovery methods, addressing the challenge of scaling analysis to larger models.
-These works collectively emphasize the importance of understanding the concrete computational steps within LLMs, moving beyond superficial observations to reveal the underlying mechanisms.
+For example, logit-lens (nostalgebraist 2020) uses the model’s decoder function to decode the intermediate activations in the vocabulary space.
+tuned-lens (Belrose et al.
+2023) refines this technique by training a different affine probe at each layer, instead of only using the pretrained model’s decoder function.
+Building on the Transformer-Lens library (Nanda and Bloom 2022), we propose Entropy-Lens, which employs logit-lens to study and characterize LLMs’ computations via their decoded version with information theory.
 Information theory in Transformers
 Information theory has been studied both in connection to the training phase of LLMs and their interpretability.
-For example, a collapse in attention entropy has been linked to training instabilities (Zhai et al., 2023) and matrix entropy was employed to evaluate “compression” in LLMs (Wei et al., 2024).
-Additionally, mutual information was used to study the effectiveness of the chain-of-thought mechanism (Ton et al., 2024).
+For example, a collapse in attention entropy has been linked to training instabilities (Zhai et al.
+2023) and matrix entropy was employed to evaluate “compression” in LLMs (Wei et al.
+2024).
+Additionally, mutual information was used to study the effectiveness of the chain-of-thought mechanism (Ton, Taufiq, and Liu 2024).
 Our work, instead, shifts the focus on the vocabulary’s natural domain.
 Through Entropy-Lens, we use information theory to study the evolution of entropy of the intermediate layers’ decoded logits.
+A related study by Dombrowski and Corlouer (2024) uses information-theoretic measures to distinguish between truthful and deceptive LLM generations.
+Their analysis focuses on deception under explicit instruction, whereas our work aims at uncovering broader entropy-based signatures across model families, prompt types, and output correctness without requiring behavioral framing or fine-tuned prompting.
 3 Background
@@
 $H(X)=-\sum_{i}p(x_{i})\log p(x_{i})=\mathbb{E}[-\log p(X)]$ (1)
-Shannon proved that this function is the only one—up to a scalar multiplication—that satisfies intuitive properties for measuring ‘disorder’.
+Shannon proved that this function is the only one—up to a scalar multiplication—that satisfies intuitive properties for measuring ‘disorder’ (Shannon 1948).
 These include being maximal for a uniform distribution, minimal for the limit of a Kronecker delta function, and ensuring that $H(A,B)\leq H(A)+H(B)$ for every possible event $A$ and $B$.
-The same function already existed in continuous form in physics, where it linked the probabilistic formalism of statistical mechanics with the more phenomenological framework of thermodynamics, where the term ‘entropy’ was originally coined.
+The same function already existed in continuous form in physics, where it linked the probabilistic formalism of statistical mechanics with the more phenomenological framework of thermodynamics, where the term ‘entropy’ was originally coined (Gibbs 1902).
+In addition to Shannon entropy, we also consider its generalization known as Rényi entropy.
+Given a discrete random variable $X$ with probability mass function $p$, the Rényi entropy of order $\alpha>0$, $\alpha\neq 1$, is defined as:
+$H_{\alpha}(X)=\frac{1}{1-\alpha}\log\sum_{i}p(x_{i})^{\alpha}$ (2)
+This formulation reduces to Shannon entropy in the limit $\alpha\to 1$, and introduces a tunable parameter $\alpha$ that modulates the sensitivity of the entropy to the distribution’s tail.
+Lower values of $\alpha$ give more weight to rare events, while higher values emphasize the most probable outcomes.
 Next, we study the entropy of vocabulary predictions—a quantity that is maximal when the prediction assigns equal probability to all tokens, minimal when it assigns zero probability to all but one token, and takes intermediate values when probability is distributed across multiple tokens, consistent with the previously mentioned properties.
+In our experiments, we explored a range of $\alpha$ values to understand how this parameter affects the informativeness of the resulting entropy profiles.
+We identified an informative range of $\alpha$ values—which includes the Shannon case $\alpha=1$—where entropy profiles retain high discriminative power for classification tasks.
+Outside of this range, we observe that profiles tend to collapse: for very small $\alpha$, entropies become nearly maximal and lose contrast across layers and examples; for large $\alpha$, they become very small and overly sensitive to noise and local fluctuations.
+These findings support the use of Shannon entropy as a balanced and parameter-free choice, robust across a wide range of practical conditions.
 3.2 The Transformer
@@
 Figure 2: Structure of a generic Transformer block.
-The transformer (Vaswani et al., 2023) is a deep learning architecture widely applied in language modelling with LLMs (Brown et al., 2020) and computer vision (Dosovitskiy et al., 2021).
+The transformer (Vaswani et al.
+2023) is a deep learning architecture widely applied in language modelling with LLMs (Brown et al.
+2020) and computer vision (Dosovitskiy et al.
+2021).
 Transformer computations happen through transformer blocks and residual connections, as exemplified in Figure 3.
-While various design choices are possible, blocks are usually a composition of layer normalization (Zhang and Sennrich, 2019), attention, and multi layer perceptrons (MLPs), as shown in Fig.
-1.
-Residual connections, instead, sum the output of the layer $i-1$ to the output of the layer $i$.
+While various design choices are possible, blocks are usually a composition of layer normalization (Zhang and Sennrich 2019), attention, and multi layer perceptrons (MLPs), as shown in Figure 2.
+Residual connections, instead, sum the output of the layer i1i-1 to the output of the layer $i$.
 Inside a single transformer block, the information flows both horizontally and vertically.
@@
 Computation schema
-LLMs are trained to simply predict the next token in a sentence.
+LLMs are trained to predict the next token in a sentence.
 That is, given a sentence prompt $S$ with tokens $t_{1},\dots,t_{N}$, the transformer encodes each token with a linear encoder $E$.
-Throughout the residual stream, the representation $x_{N}$ of the token $t_{N}$ evolves into the representation of the token $t_{N+1}$, which is then decoded back into token space via a linear decoder $D$, sometimes set to $E^{\top}$, tying the embedding weights and the decoder.
-Finally, the logits—the output of $D$—are normalized with $\operatorname{softmax}$ to represent a probability distribution over the vocabulary.
+Throughout the residual stream, the representation $x_{N}$ of the token $t_{N}$ evolves into the representation of the token $t_{N+1}$, which is then decoded back into token space via a linear decoder $D$, often set to $E^{\top}$, tying the embedding weights and the decoder.
+Finally, the logits—the output of DD—are normalized with $\operatorname{softmax}$ to represent a probability distribution over the vocabulary.
 We summarize this operation with the function $W:=\operatorname{softmax}\circ\,D$.
@@
 Hence, $x_{j}^{i}$ represents the activations of token $t_{j}$ after layer $i$.
-3.2.1 Instruct Models
-Training a Large Language Model (LLM) requires vast amounts of data and is generally divided into multiple phases.
+Instruct models
+Training an LLM requires vast amounts of data and is generally divided into multiple phases.
   • Pretraining: The model is exposed to large datasets through self-supervised tasks, such as next-token prediction or similar variants.
@@
 For our experiments, we used off-the-shelf models without RLHF to analyze information processing in a less biased transformer version.
-We also focused on Instruct models instead of Chat models for two reasons: 1.
+We also focused on Instruct models (abbreviated with it in tables and figures) instead of Chat models for two reasons: 1.
 the Instruct strategy aligns better with our experimental setup 2.
@@
 4 Method
-The aim of our framework is to find and characterize the information-theoretic signature of transformer computations. Entropy-Lens’s pipeline comprises three steps and is described in Figure 2.
+Entropy-Lens’s pipeline comprises three steps and is described in Figure 3.
+After introducing the notation and motivating the choice of entropy as a measure, in the following we describe the details of the framework which as been used for the experiments in the following sections.
 Notation
@@
 We define $W:=\operatorname{softmax}\circ\,D$ and $y_{j}^{i}:=W(x_{j}^{i})$ the normalized logits of the token $t_{j}$’s activations after layer $i$.
+Why entropy?
+Our goal is to characterize how a transformer model evolves its predictions across layers, remaining in the token space for interpretability.
+Each intermediate representation, once decoded and passed through softmax, yields a probability distribution over the vocabulary.
+However, this distribution lives on a high-dimensional, unordered categorical support.
+Classical descriptors such as variance or higher-order cumulants rely on an implicit ordering of the support and thus become meaningless when applied to token distributions—shuffling token indices alters their value arbitrarily.
+Entropy, on the other hand, is invariant under permutations and captures a well-defined notion of uncertainty or informativeness regardless of vocabulary indexing.
+It is therefore a natural and stable choice to summarize the evolution of token-level beliefs across transformer layers.
+Definitions
 The core of our methodology is to analyze the entropy of the generated tokens’ intermediate representations $y_{j}^{i}$.
@@
 where $\bigotimes$ denotes any aggregation function.
-Note that $\bigotimes$ in Definition 2 need not be the same as $\bigoplus$ defined in Definition 1.
+Note that $\bigotimes$ in Definition 2 need not be the same as $\bigoplus$ in Definition 1.
 In our experiments, we set both of them to be concatenation, so that $h_{[N:T]}$ is the matrix with $h_{N+t}$ as columns, that is $(h_{[N:T]})_{t}^{i}=h_{N+t}^{i}$ for $i\in\{1,\dots,L\}$ and $t\in\{0,\dots,T-1\}$.
@@
 The last step of our framework is classification, where we feed the aggregated entropy profile to a classifier $\mathcal{C}$ to determine whether it contains sufficient information to identify a particular ‘entity’.
-In our experiments, we examine whether aggregated entropy profiles identify model family (Section 5.1.1), task type (Section 5.1.2), and correct and wrong answers to multiple choice questions (Section 5.1.3) in LLMs.
-Additionally, we apply the same pipeline to ViTs showing the flexibility of the proposed methodology (Section 5.2).
 In our experiments, we take $\mathcal{C}$ to be a k-NN classifier.
-Classification of the aggregated entropy profiles is the step 3 our framework.
+We also examine whether aggregated entropy profiles identify model family (Section 5.1), task type and format (Sections 5.2 and 5.5), and correct and wrong answers to multiple choice questions (Section 5.3) in LLMs.
 Figure 3: Entropy-Lens’s pipeline.
@@
 5 Experiments
-The goal of our experiments is to demonstrate the effectiveness of entropy profiles in distinguishing computational signatures across various models and scenarios.
-In the first part, we explore how entropy profiles can differentiate between LLM families and characterize distinct tasks.
-In the second part, we apply our framework to the Vision Transformer architecture to assess whether insights from the LLM experiments can be similarly recovered.
-5.1 Entropy-Lens for Large Language Models
-The experiments on LLMs focus on three key aspects.
-First, we demonstrate that entropy profiles are indicative of model families, with distinctions becoming more pronounced as model size increases.
-Second, we investigate whether the entropy profile of a model alone can be used to classify the task it is performing.
-Finally, we extend this analysis to assess whether entropy profiles can distinguish between correct and incorrect task execution.
-5.1.1 Entropy profiles identify model families
-We assess whether aggregated entropy profiles can distinguish different model families by visualizing and analyzing those of 9 models from 3 different families (GPT, Gemma and LLama) with parameter counts ranging from 100M to 9B.
-We compute the mean of the entropy profiles obtained from 64 generated tokens, obtained with the prompt ‘The concept of entropy, a brief essay:’, as shown in Figure 3.
-We observe that the profiles relate uniquely to the model family, rather than a particular model, independently of its size.
+Our experiments focus on several key aspects.
+First, we show that entropy profiles are indicative of model family, with distinctions becoming more pronounced as model size increases.
+We then investigate whether a model’s entropy profile alone can be used to classify the task it is performing.
+Next, we assess whether entropy profiles capture information about output formatting, independently of task content.
+We extend this analysis to evaluate whether entropy profiles can distinguish between correct and incorrect task execution.
+Additionally, we explore the effect of varying the alpha parameter in Rényi entropy, examining how it influences the expressiveness of the resulting profiles.
+5.1 Entropy profiles identify model families
+We assess whether aggregated entropy profiles can distinguish different model families by visualizing and analyzing those of 12 models from 4 different families (GPT, Gemma, LLama and Qwen) with parameter counts ranging from 100M to 9B.
+We average 64 entropy profiles obtained by generating 32 tokens prompting the model with a completely blank prompt.
+We observe (see Figure 1) that the profiles relate uniquely to the model family, rather than a particular model, independently of its size.
+Moreover, we observe that each model size within a particular family is tied to a scaling factor if we normalize by number of layers (see Figure 4).
 The GPT model class starts with high vocabulary entropy in the early layers, indicating a wide range of possible response tokens.
 Then, entropy gradually decreases—more smoothly than in other classes—leading to a low-entropy state, where the model narrows down to a small set of possible response tokens.
-The Gemma model class, on the other hand, starts with low entropy in the very first layers, then rises to higher entropy in the intermediate layers, and finally decreases to low entropy again just before the last layers, where the model is required to produce an output token.
-The Llama model class follows a similar pattern, but with a steeper rise, resulting in a higher entropy value maintained over a larger range of intermediate layers.
+The Gemma model class, on the other hand, starts with high entropy in the very first layer, then sharply drops to lower entropy, rises again in the intermediate layers, and finally decreases to low entropy again just before the last layers, where the model is required to produce an output token.
+The Llama model class starts with low entropy , then steeply rises and maintains a high entropy value over a large range of intermediate layers, finally decreasing to low entropy again.
+The Qwen model family exhibits a similar trend, but in a more gradual manner, resulting in less well-defined regimes.
 We observe that the equivalence between models of the same family but different sizes holds when looking at the entropy trend not as a function of the absolute layer index, but rather as the relative layer position within the model.
@@
 Then, at the moment of selection, the distribution is ‘cooled down’, forcing the output to be limited to a few possible tokens.
-Figure 3: Average entropy profiles over 64 generated tokens per model. The $x$ axis is normalized for an easier comparison when models have a different number of layers.
-5.1.2 Entropy profiles identify task types
+Figure 4: Average of 64 entropy profiles over 32 generated tokens per model. The $x$ axis is normalized for an easier comparison when models have a different number of layers.
+5.2 Entropy profiles identify task types
 We verify whether the entropy profiles can identify task types examining generative (continue a text), syntactic (count the number of words in a text), and semantic (extract the subject or moral of a text) tasks.
-We do this with the TinyStories dataset (Eldan and Li, 2023).
-For evaluation robustness, we construct for each task type three prompt templates using a combination of task-specific task prompts, reported in Appendix A Table 3, and a story from TinyStories.
+We do this with the TinyStories dataset (Eldan and Li 2023).
+For evaluation robustness, we construct for each task type three prompt templates using a combination of task-specific task prompts, reported in Appendix C.1 Table 4, and a story from TinyStories.
 These templates are:
@@
 We then apply our pipeline from Section 4 to classify the aggregated entropy profiles of these prompts against their task type using a k-NN classifier.
-The model was evaluated in a 10-fold cross-validation using the ROC-AUC score (one-vs-rest).
+The model was evaluated in a 10-fold cross-validation using the ROC-AUC score (one-vs-rest), a standard choice for measuring classification performance.
 Table 1 shows the results obtained for 6 models with parameter counts ranging from 1B to 9B.
@@
 We observe that all k-NN classifiers (i.e. one for each LLM) achieve high accuracy in distinguishing entropy profiles, with a trend toward improved performance for larger models.
-Figure 4: Average entropy profiles with shaded standard deviation for different task types: generative (continuation_prompt 1 and 2), syntactic (counting_prompt 1 and 2), and semantic (semantic_prompt 1 and 2).
-These tasks are induced with the prompts described in Appendix A.
-Left: Llama-3.2-Insruct.
-Right: Gemma-2-Instruct
-Table 1: ROC-AUC (one-vs-rest) results of different models on the TinyStories task classification. The standard deviation is calculated on the 10-fold cross-validation splits.
+Figure 5: Average entropy profiles with shaded standard deviation for different task types: generative, syntactic, and semantic.
+These tasks are induced with the prompts described in Appendix C.1.
+Left: Llama-3.2-it.
+Right: Gemma-2-it.
 Model Model Size k-NN ROC-AUC
@@
 Gemma-2-it 8.9B 98.38 ± 0.50
-Llama-3.2-Instruct 1 B 94.94 ± 0.79
-Llama-3.2-Instruct 3 B 94.77 ± 0.93
-Llama-3-Instruct 8 B 96.10 ± 0.67
+Llama-3.2-it 1 B 94.94 ± 0.79
+Llama-3.2-it 3 B 94.77 ± 0.93
+Llama-3-it 8 B 96.10 ± 0.67
 Phi-3 3.6B 97.07 ± 0.87
-5.1.3 Entropy profiles identify correct task execution
-We test whether entropy profiles can identify correct and wrong answers generated by LLMs using the Massive Multitask Language Understanding (MMLU) dataset Hendrycks et al. (2021).
+Table 1: ROC-AUC (one-vs-rest) results of different models on the TinyStories task classification. Standard deviations are computed over 10-fold cross-validation splits.
+5.3 Entropy profiles identify correct task execution
+We test whether entropy profiles can identify correct and wrong answers generated by LLMs using the Massive Multitask Language Understanding (MMLU) dataset (Hendrycks et al.
+2021).
 MMLU consists of multiple-choice questions across 57 subjects, ranging from history and physics to law, mathematics, and medicine.
@@
 Each dataset entry contains: a question string, four answer choices and a label indicating the correct answer.
-We evaluated two models, a Llama-3.2 with 3B parameters Instruct and a Gemma-2 with 2B parameters, by presenting the multiple-choice questions in three different formats (as per Table 4 in Appendix A):
+We evaluate two models, a Llama-3.2 with 3B parameters Instruct and a Gemma-2 with 2B parameters, by presenting the multiple-choice questions in three different formats (as per Table 5 in Appendix C.2):
   • Base: A minimal version containing the topic, the question, and multiple-choice answers.
@@
   • Humble: A version that also instructs the model to pick a completely random option if it doesn’t know the answer.
-Then, we applied our pipeline to extract and aggregate and the responses’ entropy profiles and classify them against the correctness of the corresponding LLM-generated answer.
+Then, we apply our pipeline to extract and aggregate the responses’ entropy profiles and classify them against the correctness of the corresponding LLM-generated answer.
 We train a k-NN classifier for each LLM and validate it using 10-fold cross-validation.
-We also conducted a t-test to compare our classifier to a dummy model.
+We also conduct a t-test to compare our classifier to a dummy model.
 This dummy model generates predictions randomly, sampled from a distribution that reflects the proportion of correct and incorrect answers produced by the LLM, ensuring robustness against class imbalance.
-The results reject the null hypothesis with the k-NN achieving an AUC-ROC between 67.23 and 73.61, depending on prompt type and model (Table 2).
+The results reject the null hypothesis ($\alpha=0.05$) with the k-NN achieving an AUC-ROC between 67.23 and 73.61, depending on prompt type and model (see Table 3).
 We observe that the instruct and humble prompts improve Llama’s average accuracy, while for Gemma, this is only true for the instruct prompt.
@@
 In Gemma, however, this claim is harder to support.
-Figure 5: Entropy profiles for ViT model families.
-5.2 Entropy-Lens for Vision Transformers
-Finally, to demonstrate the versatility and robustness of our approach beyond language modeling, we analyze the entropy profiles of ViTs and DeiTs.
-Using 20 classes from ImageNet-1K (Russakovsky et al., 2015), with 20 images per class, and without any modifications to our framework, we generate the entropy profiles shown in Figure 5.
+Figure 6: PCA projection on aggregated entropy profiles extracted from the topic-format dataset.
+5.4 Three regimes of Rényi entropy
+Figure 7: Cosine similarity matrices between entropy profiles computed on a subset of the topic-format dataset using different values of $\alpha$ in the Rényi entropy.
+$\sigma$ denotes the standard deviation of the similarity matrix.
+Note how similarity flattens for very low and very high $\alpha$, while intermediate values yield more informative profiles.
+To qualitatively explore how the Rényi entropy affects the structure of entropy profiles, we compute pairwise cosine similarities between profiles generated with different values of $\alpha$.
+This analysis is performed on a subset of the topic-format dataset, and the resulting similarity matrices are shown in Figure 7.
+We observe three distinct regimes as $\alpha$ varies:
+  • For very small values of $\alpha$ (e.g., $\alpha<0.2$), the similarity matrices are nearly flat, with profiles being almost identical across examples.
+This is expected, as Rényi entropy in this regime weights all tokens with non-zero probability almost equally, and usually all tokens have non-null probability.
+  • For large values of $\alpha$ (e.g., $\alpha>20$), the similarity matrices also flatten.
+In this case, the entropy becomes increasingly dominated by the few tokens with highest probabilities.
+Since these sets of top tokens tend to have similar cardinalities (in the limit equal to 1), the profiles collapse into a narrow set of values, losing expressiveness and becoming more sensitive to local fluctuations.
+  • Between these extremes lies an informative regime—approximately $0.5\leq\alpha\leq 20$—where entropy profiles are heterogeneous enough to retain meaningful variation.
+This is reflected in the standard deviation of the similarity matrices, which peaks in this interval.
+This qualitative observation supports our empirical findings: in Section 5.5, we show that format classification accuracy remains high and stable within this informative $\alpha$ range.
+Notably, Shannon entropy ($\alpha=1$) falls within this interval, providing a strong justification for its use in the main experiments.
+By choosing $\alpha=1$, we retain discriminative power while avoiding the need to tune additional hyperparameters.
+5.5 Entropy profiles identify text format
+We test whether the entropy profiles contain signal about the format of the generated text.
+In this experiment, we prompt the model to produce short texts across different topics while enforcing one of three predefined formats: poem, scientific piece, or chat log.
+We call these generated texts the topic-format dataset.
+Model $\alpha$ k-NN ROC-AUC
+Gemma-2-2b-it 0.5 97.3 ± 1.6
+1.0 98.7 ± 1.1
+5.0 98.4 ± 1.7
+Llama-3.2-1b-it 0.5 97.8 ± 1.6
+1.0 97.8 ± 2.4
+5.0 96.6 ± 2.6
+Table 2: ROC-AUC (one-vs-rest) results of two models with three values of $\alpha$ for the Rényi entropy for format classification.
+Standard deviations are computed over 10-fold cross-validation splits.
+We use entropy profiles computed not only with Shannon entropy ($\alpha=1$), but also with Rényi entropy for $\alpha\in\{0.5,5.0\}$.
+As shown in Table 2, the k-NN classifier achieves high ROC-AUC scores across all values of $\alpha$, indicating that the format is reliably encoded in the entropy profiles.
+Moreover, the small variability in performance across $\alpha$ values supports our use of Shannon entropy as a principled, parameter-free default.
+To better understand the structure of the entropy profiles, we perform Principal Component Analysis (PCA) and visualize the first two components.
+As shown in Figure 6, the profiles cluster distinctly by format, forming linearly separable groups in the reduced space.
+This indicates that format-specific computation patterns are not only detectable by a classifier, but visibly reflected in the global shape of the entropy evolution across layers.
+Model Prompt LLM Acc. k-NN AUC-ROC
+Llama Base 50.89 73.61 ± 1.52
+Humble 58.51 69.90 ± 1.06
+Instruct 60.62 67.23 ± 1.62
+Gemma Base 56.10 71.88 ± 1.63
+Humble 54.71 72.78 ± 1.15
+Instruct 56.38 68.36 ± 1.23
+Table 3: Performance comparison of two models (Llama-3.2-3B-it and Gemma-2-2B-it) for different prompt types on MMLU accuracy and k-NN AUC-ROC score.
+6 Conclusions
+In this work, we prototyped a novel model-agnostic interpretability framework for large-scale transformer-based architectures grounded in information-theory.
+In Section 5.1, we showed that the entropy profiles of LLM intermediate predictions identify the LLM’s model family.
+Additionally, in Sections 5.2 and 5.5 we showed that the same entropy profiles can be used to identify the ‘task type’ and the ‘format type’ in LLMs, whilst in Section 5.3 we used them to distinguish between correct and wrong LLM generated answers.
+Moreover, we investigated the validity of using Shannon entropy instead of the more general Rényi entropy, and empirically demonstrated its effectiveness.
+Importantly, all our experiments were conducted on frozen off-the-shelf large-scale transformers.
+6.1 Limitations and Future Work
+While this work paves the way to further investigations in information theoretic interpretability, it also presents a number of limitations.
+First, the concepts of ‘task type’ and ‘format type’ don’t have a formal and well established definition.
+Moreover, we showed how different models possess different characteristic entropy profiles.
+We conjecture that these particular shapes are a byproduct of training procedure and architectural designs, but future research could focus on understanding the precise connections.
+Currently, we do not have a theoretical explanation for the observed phenomena or their underlying causes.
+Another interesting line of research could focus on considering more fine-grained measures of information instead of just an aggregated one such as entropy.
+With these limitations in mind, our methodology could be used to probe the reasoning capabilities of LLMs, for instance by comparing the entropy profile of a reasoning task vs.
+a data retrieval task.
+If these profiles happen to match, it could be taken as an argument against the ability of LLMs to reason.
+Conversely, if they do not, it may suggest that some form of reasoning is indeed occurring.
+Finally, recent literature explored the use of entropy for private inference (PI), where computations are performed on encrypted data without revealing users’ sensitive information (Jha and Reagen 2025).
+While previous work focused on the entropy of the attention mechanism, future research could use our methodology to develop PI-friendly applications of LLMs.
+Appendix A Appendix Structure
+The appendix is organized as follows:
+  • Appendix B – Additional Experiments: We assess the generality of our approach by applying it to Vision Transformers (ViTs).
+This analysis demonstrates that entropy profiles can be extracted and interpreted even in non-language domains, without any modification to our framework.
+  • Appendix C – Evaluation Details: We provide full details on the datasets and prompt templates used in our experiments.
+In particular, we specify how prompts were constructed for task classification, output correctness, and format classification tasks.
+This section also includes hardware information to support reproducibility.
+  • Appendix D – Minimal Implementation: We present a minimal code snippet that reproduces the core entropy profile extraction logic in a few lines of code.
+While our full codebase offers several optimizations and utilities, this section emphasizes transparency and ease of replication by showcasing the conceptual simplicity of our approach.
+Appendix B Additional Experiments
+To further test the generality and flexibility of our methodology, we conduct additional experiments beyond the core settings presented in the main text.
+In particular, we explore how Entropy-Lens performs in a different modality: computer vision.
+Without any architectural adjustment or fine-tuning, we apply the same framework to Vision Transformers (ViTs) and demonstrate that entropy profiles extracted from visual models exhibit meaningful and interpretable structure.
+This supports the claim that our method is not only scalable and model-agnostic, but also transferable across modalities.
+B.1 Entropy-Lens for Vision Transformers
+To demonstrate the versatility and robustness of our approach beyond language modeling, we analyze the entropy profiles of ViTs and DeiTs.
+Using 20 classes from ImageNet-1K (Russakovsky et al.
+2015), with 20 images per class, and without any modifications to our framework, we generate the entropy profiles shown in Figure 8.
 We observe that all profiles start with high entropy values, which then decrease, mostly in the final layers.
-This behavior is qualitatively similar to that of GPTs or larger LLaMa models (Section 5.1.1), suggesting a universal pattern across domains as different as image processing and natural language processing.
+This behavior is qualitatively similar to that of GPTs or larger LLaMa models (Section 5.1), suggesting a universal pattern across domains as different as image processing and natural language processing.
 Focusing on computer vision models, we note that while ViT and DeiT families exhibit qualitatively similar trends, they differ quantitatively—ViTs start with higher entropy values, making them easily distinguishable from DeiTs.
@@
 For ViT Large, entropy decreases more smoothly, appearing not only as a better approximation of the sharp drop seen in smaller models but possibly following a different behavior entirely, with the entropy decline starting earlier.
-We hypothesize a phase transition in entropy behavior as model size increases, occurring somewhere between $87$M and $307$M parameters.
-Table 2: Performance comparison of different models and prompt types on MMLU accuracy and k-NN AUC-ROC score.
-Model Prompt Type LLM Accuracy k-NN AUC-ROC
-Llama-3.2-3B-Instruct Base 50.89 73.61 ± 1.52
-Humble 58.51 69.90 ± 1.06
-Instruct 60.62 67.23 ± 1.62
-Gemma-2-2B-it Base 56.10 71.88 ± 1.63
-Humble 54.71 72.78 ± 1.15
-Instruct 56.38 68.36 ± 1.23
-6 Conclusions
-In this work, we prototyped a novel model-agnostic interpretability framework for large-scale transformer-based architectures grounded in information-theory.
-In Section 5.1.1, we showed that the entropy profiles of LLM intermediate predictions identify the LLM’s model family.
-In Section 5.2 we conduct similar experiments on vision transformers, demonstrating the wide applicability of our framework.
-Additionally, in Section 5.1.2 we showed that the same entropy profiles can be used to identify the ‘task type’ in LLMs, and in Section 5.1.3 we used them to distinguish between correct and wrong LLM generated answers.
-Importantly, all our experiments were conducted on frozen off-the-shelf large-scale transformers.
-Appendix A Evaluation Details
-Table 3: Prompt templates used for TinyStories tasks.
+We hypothesize a phase transition in entropy behavior as model size increases, occurring somewhere between 8787M and 307307M parameters.
+Figure 8: Entropy profiles for ViT model families.
+Appendix C Evaluation Details
+This section provides further details on the datasets and prompt templates used to evaluate the effectiveness of entropy profiles in the main experiments.
+In particular, we describe how we constructed the inputs for three key experimental settings: task type classification using the TinyStories dataset, correctness classification using the MMLU benchmark, and format classification using the topic-format dataset.
+In all cases, prompt design plays a critical role in ensuring robust comparisons across experimental conditions.
+To this end, we employed multiple prompt variations.
+The subsections below report the full set of templates used, as referenced in Sections 5.2, 5.3, and 5.5 of the main paper.
+The scripts used to generate these datasets—along with the full codebase to extract entropy profiles and reproduce all experiments—are shared as part of our code release.
+Finally, we also provide information about the hardware used to run our experiments to facilitate reproducibility.
+C.1 Prompt templates for TinyStories tasks
+In Section 5.2, we describe an experimental setup designed to test whether entropy profiles can identify different types of tasks.
+To this end, we used the TinyStories dataset (Eldan and Li 2023) and constructed prompts combining short stories with specific task instructions.
+Each task type—generative, syntactic, and semantic—was associated with two distinct natural language formulations, referred to as task prompts.
+These are listed in Table 4.
+By varying the task prompt, we ensure that our classification results are not simply driven by surface-level textual artifacts, but instead reflect deeper computational signatures captured by the entropy profiles.
+This table complements the prompt templates (base, reversed, and scrambled) described in the main text and provides the full set of instructions used to elicit different model behaviors.
 Task Type Task prompt
@@
 Who is the subject of the story?
-Table 4: Prompt templates used for the MMLU dataset.
+Table 4: Prompt templates used for TinyStories tasks.
+C.2 Prompt templates for MMLU correctness classification
+In Section 5.3, we test whether entropy profiles can distinguish between correct and incorrect answers produced by language models.
+To construct the dataset for this experiment, we used the Massive Multitask Language Understanding (MMLU) benchmark (Hendrycks et al.
+2021), applying three distinct prompt styles to elicit different answer behaviors from the models.
+Table 5 reports the full prompt templates used in this experiment.
+Each template presents the same multiple-choice question in a different instructional format: the Base format presents the question directly; the Instruct format introduces an explicit instruction to select a single correct answer; and the Humble format includes an additional fallback directive encouraging the model to guess randomly if uncertain.
+This variation in prompting allows us to control for instruction framing and to evaluate whether entropy profiles can capture response confidence and correctness robustly across different model behaviors.
+The table shown here complements the description in the main text.
 Prompt Type Prompt
@@
 Answer:
+Table 5: Prompt templates used for the MMLU dataset.
+C.3 Prompt Construction for the Topic-Format Dataset
+To evaluate whether entropy profiles captures stylistic features of generated text, we constructed a custom dataset referred to as the topic-format dataset.
+In this setting, models are prompted to generate short texts on various topics, each constrained to adopt one of three specific formats: poem, scientific piece, or chat log.
+The goal is to determine whether these formats induce distinct entropy profiles, independently of the topic content.
+We generated prompts by pairing 150 distinct topics with the following three format instructions:
+  • Write a poem about ...
+  • Write a scientific piece about ...
+  • Simulate a chat log about ...
+Each prompt is constructed by concatenating a format prefix with a randomly selected topic (e.g., Write a poem about a planet).
+The resulting dataset contains 450 prompt completions per model, each paired with its entropy profile computed using Rényi entropy for $\alpha\in\{0.5,1.0,5.0\}$.
+All generations were performed using a maximum generation length of 256 tokens.
+We then segmented the output into 8 equal-length windows and computed an entropy profile for each.
+The resulting data were stored with format labels and used in the classification and visualization tasks discussed in Section 5.5 of the main paper.
+This setup enables robust testing of the extent to which entropy profiles encode formatting cues, beyond topical content or task semantics.
+C.4 Experimental and hardware setup
+All experiments were conducted on a compute node equipped with an NVIDIA L40 GPU, an Intel Xeon Gold CPU, 128 GB of RAM, and running Ubuntu 22.04.
+The primary software frameworks used were PyTorch, Transformer-Lens, and HuggingFace Transformers.
+Inference on LLMs was performed using float16 precision for improved efficiency.
+Appendix D Minimal Implementation
+To maximize reproducibility and transparency, we provide a minimal implementation of our framework.
+While our full codebase includes optimizations and utility functions to streamline analysis across models and datasets, the core idea behind Entropy-Lens is conceptually simple and can be expressed in just a few lines of code.
+This section presents a compact example that computes the entropy profile of generated tokens using an off-the-shelf language model.
+Despite its brevity, this snippet captures the essence of our method: extracting intermediate representations, mapping them to vocabulary distributions, and computing their entropies.
+D.1 Minimal Entropy Profile Extraction
+The code in listing 1 demonstrates how to compute an entropy profile for a single prompt using a standard decoder-only transformer.
+It relies only on model forward passes and the use of logit-lens-style decoding.
+No gradients, fine-tuning, or access to internal model weights are required.
+1from transformers import AutoTokenizer, AutoModelForCausalLM
+2import torch
+3
+4# Load GPT-2 and set up
+5tokenizer = AutoTokenizer.from_pretrained(’gpt2’)
+6model = AutoModelForCausalLM.from_pretrained(’gpt2’, device_map="auto").eval()
+7tokenizer.pad_token = tokenizer.eos_token
+8
+9# Define entropy computation
+10ln, U = model.transformer.ln_f, model.lm_head
+11entropy = lambda x: -torch.sum(x * torch.log(x + 1e-15), dim=-1)
+12
+13# Prepare input
+14input_text = ’The concept of entropy’
+15inputs = tokenizer.encode(input_text, return_tensors="pt").to(model.device)
+16
+17# Generate with hidden states
+18outputs = model.generate(inputs,
+19 do_sample=True,
+20 max_new_tokens=32,
+21 output_hidden_states=True,
+22 return_dict_in_generate=True,
+23 pad_token_id=tokenizer.pad_token_id
+24)
+25
+26# Stack hidden activations and compute entropy signature
+27activations = torch.vstack([
+28 torch.vstack(h).permute(1, 0, 2) for h in outputs.hidden_states
+29])
+30entropy_signature = entropy(U(ln(activations)).softmax(dim=-1))
+Listing 1: A minimal Python implementation of Entropy-Lens using Huggingface models and Pytorch.

### v3

@@
-Entropy-Lens: The Information Signature of Transformer Computations
+Entropy-Lens: Uncovering Decision Strategies in LLMs
 Abstract
-Transformer models map input token sequences to output token distributions, layer by layer.
-While most interpretability work focuses on internal latent representations, we study the evolution of these token-level distributions directly in vocabulary space.
-However, such distributions are high-dimensional and defined on an unordered support, making common descriptors like moments or cumulants ill-suited.
-We address this by computing the Shannon entropy of each intermediate predicted distribution, yielding one interpretable scalar per layer.
-The resulting sequence—the entropy profile—serves as a compact, information-theoretic signature of the model’s computation.
-We introduce Entropy-Lens, a model-agnostic framework that extracts entropy profiles from frozen, off-the-shelf transformers.
-We show that these profiles (i) reveal family-specific computation patterns invariant under depth rescaling, (ii) are predictive of prompt type and task format, and (iii) correlate with output correctness.
-We further show that Rényi entropies yield similar results within a broad range of $\alpha$ values, justifying the use of Shannon entropy as a stable and principled summary.
-Our results hold across different transformers, without requiring gradients, fine-tuning, or access to model internals.
-00footnotetext: Preprint - This manuscript is currently under review.11footnotetext: Code available at: github.com/christopher-irw/Entropy-Lens
+In large language models (LLMs), each block operates on the residual stream to map input token sequences to output token distributions.
+However, most of the interpretability literature focuses on internal latent representations, leaving token-space dynamics underexplored.
+The high dimensionality and categoricity of token distributions hinder their analysis, as standard statistical descriptors are not suitable.
+We show that the entropy of logit-lens predictions overcomes these issues.
+In doing so, it provides a per-layer scalar, permutation-invariant metric.
+We introduce Entropy-Lens to distill the token-space dynamics of the residual stream into a low-dimensional signal.
+We call this signal the entropy profile.
+We apply our method to a variety of model sizes and families, showing that (i) entropy profiles uncover token prediction dynamics driven by expansion and pruning strategies; (ii) these dynamics are family-specific and invariant under depth rescaling; (iii) they are characteristic of task type and output format; (iv) these strategies have unequal impact on downstream performance, with the expansion strategy usually being more critical.
+Ultimately, our findings further enhance our understanding of the residual stream, enabling a granular assessment of how information is processed across model depth.
+Machine Learning, ICML
 1 Introduction
-Transformer-based architectures (Vaswani et al.
-2023) operate by iteratively mapping token inputs to distributions over next-token outputs, layer by layer (Shan et al.
-2024).
-This view emphasizes their inherently probabilistic nature: each layer produces a distribution over the vocabulary, encoding the model’s beliefs at that stage.
-Despite their success across domains—from language to biology to vision (Devlin 2018; Ji et al.
-2021; Wu et al.
-2020)—the internal evolution of these distributions remains poorly understood, resulting in unpredictable behaviour (Wei et al.
-2022) and reliability concerns (Schroeder and Wood-Doughty 2025; Huang et al.
-2025).
-Most interpretability research either focuses on toy models (Elhage et al.
-2021, 2022) or requires training a set of probes (nostalgebraist 2020; Belrose et al.
-2023) or full models on ad-hoc tasks (Nanda et al.
-2023), limiting scalability and generality.
-Moreover, existing methods primarily study latent representations—hidden states, attention maps, or MLP activations (Nanda and Bloom 2022; Bereska and Gavves 2024; Chefer, Gur, and Wolf 2021)—rather than the output distributions themselves.
-Figure 1: TSNE of the aggregated entropy profiles of different model families using unprompted models. The size of each point is scaled based on the number of parameters of the model.
-We propose to shift perspective: instead of analyzing the internal geometry of latent activations, we focus on the evolution of the model’s token-level predictions across layers.
-These predictions, obtained by decoding intermediate representations via the model’s output head, form probability distributions over the vocabulary.
-Studying their evolution offers a direct, vocabulary-grounded view of the computation.
-However, these distributions pose two challenges: they are high-dimensional (one value per token) and defined over an unordered support (the vocabulary).
-This makes classical descriptors like variance or higher-order cumulants unstable or ill-defined.
-We address both issues by computing the Shannon entropy of each distribution—a scalar, interpretable quantity invariant to token permutation, and reflective of the model’s uncertainty at each layer.
-We introduce Entropy-Lens, a simple, scalable, and model-agnostic framework for studying transformer computations through the evolution of entropy across layers.
-In our experiments, we consider several LLMs, including Llama (Touvron et al.
-2023), Gemma (Team et al.
-2024), and GPT, (Radford 2018) up to 9B parameters.
-For each generated token, we compute the entropy of its intermediate predictions, as described in Section 4, yelding one interpretable value per layer: this constitutes the entropy profile.
-We then aggregate these profiles across tokens and study whether they capture structural information about the internal LLM’s computation (Figure 1).
-In Section 5, we show that: (i) entropy profiles reveal family-specific signatures of transformer architectures, invariant under depth rescaling; (ii) they are predictive of prompt semantics and format; (iii) they correlate with output correctness; (iv) and results are stable across Rényi entropy variants, justifying our use of Shannon entropy as a principled default.
-Entropy-Lens requires no gradients, no fine-tuning, and no access to internal weights.
-It applies to frozen, off-the-shelf transformers of arbitrary size.
-While our experiments focus on language models, the methodology is general and opens new directions for vocabulary-grounded interpretability.
-2 Related Work
-Lenses in LLMs
-Mechanistic interpretability (Bereska and Gavves 2024) aims to provide a precise description and prediction of transformer-based computations.
-Common tools in the field are lenses, which are a broad class of probes deployed in intermediate steps of the residual stream.
-For example, logit-lens (nostalgebraist 2020) uses the model’s decoder function to decode the intermediate activations in the vocabulary space.
-tuned-lens (Belrose et al.
-2023) refines this technique by training a different affine probe at each layer, instead of only using the pretrained model’s decoder function.
-Building on the Transformer-Lens library (Nanda and Bloom 2022), we propose Entropy-Lens, which employs logit-lens to study and characterize LLMs’ computations via their decoded version with information theory.
-Information theory in Transformers
-Information theory has been studied both in connection to the training phase of LLMs and their interpretability.
-For example, a collapse in attention entropy has been linked to training instabilities (Zhai et al.
-2023) and matrix entropy was employed to evaluate “compression” in LLMs (Wei et al.
-2024).
-Additionally, mutual information was used to study the effectiveness of the chain-of-thought mechanism (Ton, Taufiq, and Liu 2024).
-Our work, instead, shifts the focus on the vocabulary’s natural domain.
-Through Entropy-Lens, we use information theory to study the evolution of entropy of the intermediate layers’ decoded logits.
-A related study by Dombrowski and Corlouer (2024) uses information-theoretic measures to distinguish between truthful and deceptive LLM generations.
-Their analysis focuses on deception under explicit instruction, whereas our work aims at uncovering broader entropy-based signatures across model families, prompt types, and output correctness without requiring behavioral framing or fine-tuned prompting.
+Transformer-based architectures (Vaswani et al., 2023) consist of stacked blocks.
+Each one operates on the residual stream by iteratively mapping input tokens to output distributions over the next-token (Shan et al., 2024).
+Much of the interpretability literature studies the residual stream in embedding space (Skean et al., 2025; Jawahar et al., 2019; Elhage et al., 2022), often by analyzing geometric arrangements and dynamics of aggregated latent representations (Skean et al., 2025; Sarfati et al., 2025; Li et al., 2025).
+These methods offer useful diagnostics but fail to offer a low-dimensional, interpretable, and per-sample representation of the model’s behavior in token space (Sarfati et al., 2025).
+We tackle this gap by projecting the residual stream into the token space via logit-lens (nostalgebraist, 2020) and analyzing the entropy dynamics of the decoded distributions.
+With this approach, grounded in information theory, we distill the residual stream into a low-dimensional signal.
+This allows us to track token prediction dynamics layer-by-layer and on individual generations.
+Our framework, Entropy-Lens, uses entropy as a natural metric to uncover token prediction dynamics driven by two fundamental strategies: (1) expansion, where the model considers new candidates for the next token, and (2) pruning, where it refines the candidate set.
+We compute the entropy of the logit-lens probabilities as a count of the number of possible candidates for each layer and sample.
+Our analysis reveals that model families have a characteristic way of mixing these two strategies across layers, which is consistent between models of different sizes in the same family.
+Furthermore, we show that different tasks (e.g.
+generative, syntactic, semantic) and output-formats (e.g.
+poem, scientific piece, chat log) induce distinct token prediction dynamics.
+Finally, we show in a multiple-choice setup that expansion and pruning strategies have different relative importance for downstream performance.
+At the same time, we observe an overall trend emphasizing the expansion phases.
+In summary, our main contributions are:
+  • We propose entropy as a sound metric to study the token prediction dynamics and expansion/pruning strategies in the transformer body with single layer and single token level granularity.
+  • We show that the particular mixing between these strategies is characteristic of model families, task type and output format.
+  • We show that, given a particular model, expansion/pruning strategies have different relative importance for downstream performance, with an overall trend emphasizing expansion.
 2 Background
+In this section, we review the fundamental concepts from information theory and key components of the Transformer architecture, with particular emphasis on the residual connections and the information flow across layers.
+We also briefly categorize the training stages of Large Language Models to clarify the distinction between the base and instruct models used in our study.
 2.1 Information theory
+Shannon entropy.
 The main information-theoretic quantity used in our study is entropy.
@@
 random variable $X$ with outcomes $x_{i}$ and probability mass function $p$, the Shannon entropy $H$ of $X$ is defined as
-$H(X)=-\sum_{i}p(x_{i})\log p(x_{i})=\mathbb{E}[-\log p(X)]$ (1)
-Shannon proved that this function is the only one—up to a scalar multiplication—that satisfies intuitive properties for measuring ‘disorder’ (Shannon 1948).
-These include being maximal for a uniform distribution, minimal for the limit of a Kronecker delta function, and ensuring that $H(A,B)\leq H(A)+H(B)$ for every possible event $A$ and $B$.
-The same function already existed in continuous form in physics, where it linked the probabilistic formalism of statistical mechanics with the more phenomenological framework of thermodynamics, where the term ‘entropy’ was originally coined (Gibbs 1902).
+$H(X)=-\sum_{i}p(x_{i})\log p(x_{i})=\mathbb{E}[-\log p(X)].$ (1)
+Shannon proved that this function is the only one, up to a scalar multiplication, which satisfies intuitive properties for measuring ‘disorder’ (Shannon, 1948).
+These include being maximal for a uniform distribution, minimal for the limit of a Kronecker delta function, and ensuring that $H(A,B)\leq H(A)+H(B)$ for every possible random variable $A$ and $B$.
+The same function already existed in continuous form in physics, where it linked the probabilistic formalism of statistical mechanics with the more phenomenological framework of thermodynamics, where the term ‘entropy’ was originally coined (Gibbs, 1902).
+Rényi entropy.
 In addition to Shannon entropy, we also consider its generalization known as Rényi entropy.
 Given a discrete random variable $X$ with probability mass function $p$, the Rényi entropy of order $\alpha>0$, $\alpha\neq 1$, is defined as:
-$H_{\alpha}(X)=\frac{1}{1-\alpha}\log\sum_{i}p(x_{i})^{\alpha}$ (2)
+$H_{\alpha}(X)=\frac{1}{1-\alpha}\log\sum_{i}p(x_{i})^{\alpha}.$ (2)
 This formulation reduces to Shannon entropy in the limit $\alpha\to 1$, and introduces a tunable parameter $\alpha$ that modulates the sensitivity of the entropy to the distribution’s tail.
-Lower values of $\alpha$ give more weight to rare events, while higher values emphasize the most probable outcomes.
-Next, we study the entropy of vocabulary predictions—a quantity that is maximal when the prediction assigns equal probability to all tokens, minimal when it assigns zero probability to all but one token, and takes intermediate values when probability is distributed across multiple tokens, consistent with the previously mentioned properties.
-In our experiments, we explored a range of $\alpha$ values to understand how this parameter affects the informativeness of the resulting entropy profiles.
-We identified an informative range of $\alpha$ values—which includes the Shannon case $\alpha=1$—where entropy profiles retain high discriminative power for classification tasks.
-Outside of this range, we observe that profiles tend to collapse: for very small $\alpha$, entropies become nearly maximal and lose contrast across layers and examples; for large $\alpha$, they become very small and overly sensitive to noise and local fluctuations.
-These findings support the use of Shannon entropy as a balanced and parameter-free choice, robust across a wide range of practical conditions.
+Rényi entropy also subsumes many classical descriptors of discrete distributions without intrinsic ordering: with appropriate choices of $\alpha$, it recovers collision entropy ($\alpha=2$), min-entropy ($\alpha\to\infty$), and max-entropy ($\alpha\to 0$), and it correlates with indices such as the Gini–Simpson index (Rényi, 1961; Jost, 2006).
+In general, lower values of $\alpha$ give more weight to rare events, while higher values emphasize the most probable outcomes.
+Top-pp.
+Also referred to as Nucleus Sampling (Holtzman et al., 2020), Top-$p$ is a thresholding technique that restricts the distribution to the smallest set of tokens whose cumulative probability mass exceeds a value $p\in(0,1]$.
 2.2 The Transformer
-Architecture
-Figure 2: Structure of a generic Transformer block.
-The transformer (Vaswani et al.
-2023) is a deep learning architecture widely applied in language modelling with LLMs (Brown et al.
-2020) and computer vision (Dosovitskiy et al.
-2021).
+Architecture.
+The transformer (Vaswani et al., 2023) is a deep learning architecture widely applied in language modelling with LLMs (Brown et al., 2020) and computer vision (Dosovitskiy et al., 2021).
 Transformer computations happen through transformer blocks and residual connections, as exemplified in Figure 2.
-While various design choices are possible, blocks are usually a composition of layer normalization (Zhang and Sennrich 2019), attention, and multi layer perceptrons (MLPs), as shown in Figure 2.
-Residual connections, instead, sum the output of the layer i1i-1 to the output of the layer $i$.
+While various design choices are possible, blocks are usually a composition of layer normalization (Zhang and Sennrich, 2019), attention, and multi layer perceptrons (MLPs).
+Residual connections, instead, sum the output of layer i1i-1 to the output of layer $i$.
 Inside a single transformer block, the information flows both horizontally and vertically.
 The former, enabled by the attention mechanism, allows the token representations to interact with each other.
-In a language modelling task, for example, this is useful to identify which parts of the input sequence—the sentence prompt—should influence the next token prediction and quantify by how much.
+In a language modelling task, for example, this is useful to identify which parts of the input sequence, the sentence prompt, should influence the next token prediction and quantify by how much.
 The latter vertical information flow allows the representation to evolve and encode different meanings or concepts.
@@
 The embedding spaces where these computations take place are generally called the residual stream.
-Computation schema
+Computation schema.
 LLMs are trained to predict the next token in a sentence.
 That is, given a sentence prompt $S$ with tokens $t_{1},\dots,t_{N}$, the transformer encodes each token with a linear encoder $E$.
-Throughout the residual stream, the representation $x_{N}$ of the token $t_{N}$ evolves into the representation of the token $t_{N+1}$, which is then decoded back into token space via a linear decoder $D$, often set to $E^{\top}$, tying the embedding weights and the decoder.
-Finally, the logits—the output of DD—are normalized with $\operatorname{softmax}$ to represent a probability distribution over the vocabulary.
-We summarize this operation with the function $W:=\operatorname{softmax}\circ\,D$.
-In formal terms, information processing can be expressed using the encoder, decoder, Transformer block $f$, and residual connection:
-$x_{j}^{0}$ $=E(t_{j})$ (3)
-$x_{j}^{i}$ $=f^{i}(x_{j}^{i-1})+x_{j}^{i-1}$ (4)
-$y_{j}^{i}$ $=W(x_{j}^{i})$ (5)
-where $j\in\{1,\dots,N\}$ ranges over the number of tokens in the prompt and $i\in\{0,\dots,L\}$ ranges over the number of layers.
-Hence, $x_{j}^{i}$ represents the activations of token $t_{j}$ after layer $i$.
-Instruct models
-Training an LLM requires vast amounts of data and is generally divided into multiple phases.
-  • Pretraining: The model is exposed to large datasets through self-supervised tasks, such as next-token prediction or similar variants.
-This phase helps the model learn a broad range of general knowledge.
-  • Fine-tuning: This phase teaches the model to generate more useful and coherent responses.
-Two main strategies are used: Chat: The model is trained on structured conversations between a user and the model, with clearly defined roles.
-Instruct: The model learns from simple commands, without a predefined dialogue structure.
-  • RLHF (optional): Some models undergo Reinforcement Learning from Human Feedback (RLHF) to further refine their responses based on human preferences.
-For our experiments, we used off-the-shelf models without RLHF to analyze information processing in a less biased transformer version.
-We also focused on Instruct models (abbreviated with it in tables and figures) instead of Chat models for two reasons: 1.
-the Instruct strategy aligns better with our experimental setup 2.
-Instruct models are more flexible and often preferred for practical applications.
-4 Method
-Entropy-Lens’s pipeline comprises three steps and is described in Figure 3.
-After introducing the notation and motivating the choice of entropy as a measure, in the following we describe the details of the framework which as been used for the experiments in the following sections.
-Notation
-We denote the input sentence comprising tokens $t_{1},\dots,t_{N}$ by $S=(t_{i})_{i=1}^{N}$.
-Then, $x_{j}^{i}$ denotes the activations of the token $t_{j}$ after block $i$ for $j\in\{1,\dots,N\}$ and $i\in\{1,\dots,L\}$.
-Since our analysis focuses on the logits extracted from the intermediate layers of the transformer, it will be useful to distinguish between normalized and unnormalized logits.
-We define $W:=\operatorname{softmax}\circ\,D$ and $y_{j}^{i}:=W(x_{j}^{i})$ the normalized logits of the token $t_{j}$’s activations after layer $i$.
-Why entropy?
-Our goal is to characterize how a transformer model evolves its predictions across layers, remaining in the token space for interpretability.
-Each intermediate representation, once decoded and passed through softmax, yields a probability distribution over the vocabulary.
-However, this distribution lives on a high-dimensional, unordered categorical support.
-Classical descriptors such as variance or higher-order cumulants rely on an implicit ordering of the support and thus become meaningless when applied to token distributions—shuffling token indices alters their value arbitrarily.
-Entropy, on the other hand, is invariant under permutations and captures a well-defined notion of uncertainty or informativeness regardless of vocabulary indexing.
-It is therefore a natural and stable choice to summarize the evolution of token-level beliefs across transformer layers.
-Definitions
-The core of our methodology is to analyze the entropy of the generated tokens’ intermediate representations $y_{j}^{i}$.
-These vectors are probability distributions, as they are the output of a $\operatorname{softmax}$.
-To obtain a single quantity that summarizes the information they contain, we compute their entropy $H(y_{j}^{i})$.
-For one generated token, we can consider the entropy of all of its intermediate predictions $H(y_{j}^{i})$ for $i\in\{1,\dots,L\}$.
-This leads us to the definition of entropy profile:
-Defintion 1 (Entropy profile)
-Let $h_{j}^{i}=H(y_{j}^{i})$ be the entropy of the intermediate representation of token $t_{j}$ after block $i$ and residual connection. The entropy profile of the next generated token is defined as
-$h_{N}=\bigoplus_{i}h_{N}^{i}$ (6)
-where $\bigoplus$ denotes any aggregation function.
-In our experiments, we set $\bigoplus$ to be concatenation, so that $h_{N}=(h_{N}^{1},\dots,h_{N}^{L})^{\top}$, but other choices are possible.
-The extraction of entropy profiles is the step 1 of our pipeline.
-Then, we fix the number of tokens that the LLM is required to generate, $T$ and repeat the same procedure for each of them, leading us to the next definition:
-Defintion 2 (Aggregated entropy profile)
-Let $h_{N+t}$ be the entropy profiles according to Definition 1 for $t\in\{0,\dots,T-1\}$, i.e.
-the entropy profile of each token generated sequentially by a transformer.
-The aggregated entropy profile of the next $T$ generated tokens is defined as
-$h_{[N:T]}=\bigotimes_{t=0}^{T-1}h_{N+t}$ (7)
-where $\bigotimes$ denotes any aggregation function.
-Note that $\bigotimes$ in Definition 2 need not be the same as $\bigoplus$ in Definition 1.
-In our experiments, we set both of them to be concatenation, so that $h_{[N:T]}$ is the matrix with $h_{N+t}$ as columns, that is $(h_{[N:T]})_{t}^{i}=h_{N+t}^{i}$ for $i\in\{1,\dots,L\}$ and $t\in\{0,\dots,T-1\}$.
-The aggregation of entropy profiles is the step 2 of our framework.
-The last step of our framework is classification, where we feed the aggregated entropy profile to a classifier $\mathcal{C}$ to determine whether it contains sufficient information to identify a particular ‘entity’.
-In our experiments, we take $\mathcal{C}$ to be a k-NN classifier.
-We also examine whether aggregated entropy profiles identify model family (Section 5.1), task type and format (Sections 5.2 and 5.5), and correct and wrong answers to multiple choice questions (Section 5.3) in LLMs.
-Figure 3: Entropy-Lens’s pipeline.
-(Left) A diagram representing a transformer architecture: hidden representations are converted into intermediate predictions with $W$ before calculating their entropy $H$.
-(Right) A diagram representing our framework: step 1: entropy profile extraction, step 2: entropy profile aggregation and step 3: classification.
-5 Experiments
-Our experiments focus on several key aspects.
-First, we show that entropy profiles are indicative of model family, with distinctions becoming more pronounced as model size increases.
-We then investigate whether a model’s entropy profile alone can be used to classify the task it is performing.
-Next, we assess whether entropy profiles capture information about output formatting, independently of task content.
-We extend this analysis to evaluate whether entropy profiles can distinguish between correct and incorrect task execution.
-Additionally, we explore the effect of varying the alpha parameter in Rényi entropy, examining how it influences the expressiveness of the resulting profiles.
-5.1 Entropy profiles identify model families
-We assess whether aggregated entropy profiles can distinguish different model families by visualizing and analyzing those of 12 models from 4 different families (GPT, Gemma, LLama and Qwen) with parameter counts ranging from 100M to 9B.
-We average 64 entropy profiles obtained by generating 32 tokens prompting the model with a completely blank prompt.
-We observe (see Figure 1) that the profiles relate uniquely to the model family, rather than a particular model, independently of its size.
-Moreover, we observe that each model size within a particular family is tied to a scaling factor if we normalize by number of layers (see Figure 4).
-The GPT model class starts with high vocabulary entropy in the early layers, indicating a wide range of possible response tokens.
-Then, entropy gradually decreases—more smoothly than in other classes—leading to a low-entropy state, where the model narrows down to a small set of possible response tokens.
-The Gemma model class, on the other hand, starts with high entropy in the very first layer, then sharply drops to lower entropy, rises again in the intermediate layers, and finally decreases to low entropy again just before the last layers, where the model is required to produce an output token.
-The Llama model class starts with low entropy , then steeply rises and maintains a high entropy value over a large range of intermediate layers, finally decreasing to low entropy again.
-The Qwen model family exhibits a similar trend, but in a more gradual manner, resulting in less well-defined regimes.
-We observe that the equivalence between models of the same family but different sizes holds when looking at the entropy trend not as a function of the absolute layer index, but rather as the relative layer position within the model.
-We conjecture that high entropy phases, whether in the early or intermediate layers, allow the model to explore more possibilities in its response, similarly to how temperature helps avoid getting stuck in local minima in optimization.
-Then, at the moment of selection, the distribution is ‘cooled down’, forcing the output to be limited to a few possible tokens.
-Figure 4: Average of 64 entropy profiles over 32 generated tokens per model. The $x$ axis is normalized for an easier comparison when models have a different number of layers.
-5.2 Entropy profiles identify task types
-We verify whether the entropy profiles can identify task types examining generative (continue a text), syntactic (count the number of words in a text), and semantic (extract the subject or moral of a text) tasks.
-We do this with the TinyStories dataset (Eldan and Li 2023).
-For evaluation robustness, we construct for each task type three prompt templates using a combination of task-specific task prompts, reported in Appendix C.1 Table 4, and a story from TinyStories.
-These templates are:
-  • Base, of the form task prompt + story
-  • Reversed, of the form story + task prompt
-  • Scrambled, of the form task prompt + scrambled story or scrambled story + task prompt, at random.
-A scrambled story is a ‘story’ obtained by randomly shuffling the words in a given story from TinyStories.
-Note that, for a robust evaluation, we also use 2 possible task prompt variations, as per Table 4.
-We generate 800 prompts per task type, $1/3$ of them with the base template, $1/3$ with the reversed template, and $1/3$ with the scrambled templates, for a total of 2400 prompts.
-We then apply our pipeline from Section 4 to classify the aggregated entropy profiles of these prompts against their task type using a k-NN classifier.
-The model was evaluated in a 10-fold cross-validation using the ROC-AUC score (one-vs-rest), a standard choice for measuring classification performance.
-Table 1 shows the results obtained for 6 models with parameter counts ranging from 1B to 9B.
-Figure 5 shows the average entropy profiles per task type.
-We observe that all k-NN classifiers (i.e. one for each LLM) achieve high accuracy in distinguishing entropy profiles, with a trend toward improved performance for larger models.
-Figure 5: Average entropy profiles with shaded standard deviation for different task types: generative, syntactic, and semantic.
-These tasks are induced with the prompts described in Appendix C.1.
-Left: Llama-3.2-it.
-Right: Gemma-2-it.
-Model Model Size k-NN ROC-AUC
-Gemma-2-it 2.1B 97.66 ± 0.47
-Gemma-2-it 8.9B 98.38 ± 0.50
-Llama-3.2-it 1 B 94.94 ± 0.79
-Llama-3.2-it 3 B 94.77 ± 0.93
-Llama-3-it 8 B 96.10 ± 0.67
-Phi-3 3.6B 97.07 ± 0.87
-Table 1: ROC-AUC (one-vs-rest) results of different models on the TinyStories task classification. Standard deviations are computed over 10-fold cross-validation splits.
-5.3 Entropy profiles identify correct task execution
-We test whether entropy profiles can identify correct and wrong answers generated by LLMs using the Massive Multitask Language Understanding (MMLU) dataset (Hendrycks et al.
-2021).
-MMLU consists of multiple-choice questions across 57 subjects, ranging from history and physics to law, mathematics, and medicine.
-The difficulty levels span from elementary to professional, making it a benchmark for evaluating language models on specialized domains.
-Each dataset entry contains: a question string, four answer choices and a label indicating the correct answer.
-We evaluate two models, a Llama-3.2 with 3B parameters Instruct and a Gemma-2 with 2B parameters, by presenting the multiple-choice questions in three different formats (as per Table 5 in Appendix C.2):
-  • Base: A minimal version containing the topic, the question, and multiple-choice answers.
-  • Instruct: A version with a brief explanation that it’s a multiple-choice test where only one option should be selected.
-  • Humble: A version that also instructs the model to pick a completely random option if it doesn’t know the answer.
-Then, we apply our pipeline to extract and aggregate the responses’ entropy profiles and classify them against the correctness of the corresponding LLM-generated answer.
-We train a k-NN classifier for each LLM and validate it using 10-fold cross-validation.
-We also conduct a t-test to compare our classifier to a dummy model.
-This dummy model generates predictions randomly, sampled from a distribution that reflects the proportion of correct and incorrect answers produced by the LLM, ensuring robustness against class imbalance.
-The results reject the null hypothesis ($\alpha=0.05$) with the k-NN achieving an AUC-ROC between 67.23 and 73.61, depending on prompt type and model (see Table 3).
-We observe that the instruct and humble prompts improve Llama’s average accuracy, while for Gemma, this is only true for the instruct prompt.
-Additionally, in Llama, the model’s higher accuracy seems to be partially linked to greater difficulty in distinguishing correct from incorrect entropy profiles, though more rigorous analysis is needed to confirm this.
-In Gemma, however, this claim is harder to support.
-Figure 6: PCA projection on aggregated entropy profiles extracted from the topic-format dataset.
+Throughout the residual stream, the representation $\mathbf{x}_{N}$ of the token $t_{N}$ evolves into the representation of the token $t_{N+1}$, which is then decoded back into token space via a linear decoder $D$, often set to $E^{\top}$, tying the embedding weights and the decoder.
+Finally, the logits, the output of $D$, are normalized with $\operatorname{softmax}$ to represent a probability distribution over the vocabulary.
+Instruct models.
+Training an LLM typically progresses from self-supervised pretraining on vast corpora to fine-tuning on structured instructions, often followed by RLHF for alignment (Ouyang et al., 2022).
+While distinction varies, Chat tuning focuses on dialogue history, whereas Instruct tuning (Wei et al., 2022) targets adherence to specific commands.
+In this work, we utilize off-the-shelf Instruct models (denoted ‘it’).
+We prioritize these over specific chat variants as their flexibility in following command-style prompts aligns better with our experimental setup.
+3 Entropy and decision strategies
+In this section, we motivate the use of entropy from two perspectives.
+First, Section 3.1 establishes it as a robust dimensionality reduction technique for high-dimensional, unordered vocabulary distributions.
+This addresses the main challenges posed by high-dimensional unordered distributions over token vocabularies.
+Second, Section 3.2 argue that changes in entropy $\Delta H_{i}$ between a layer $i$ and its predecessor provide a mechanistic interpretation, serving as indicators of expansion and pruning of the next-token candidate set.
+Finally, Section 3.3 generalizes our framework to the broader family of Rényi entropies.
+In particular, we argue that our findings are robust across a wide and informative range of $\alpha$ values, and not a simple arbitrary choice.
+3.1 Entropy as dimensionality reduction
+Analyzing distributions produced by transformers poses two challenges.
+First, they are high-dimensional, with one probability per vocabulary token.
+Second, there is no intrinsic notion of order between tokens in the vocabulary.
+As a result, many standard statistical summaries such as moments or cumulants are unstable or ill-defined when applied to vocabulary distributions.
+We address both issues by computing the entropy of each distribution, which has two properties: (i) it is a scalar quantity, and (ii) it is invariant to token permutation.
+Property (i) addresses the high-dimensionality challenge, and property (ii) makes it suitable for unordered categorical distributions.
+In the main experiments, we use Shannon entropy $H$, defined in Section 2.
+Section 3.3 discusses the broader class of Rényi entropies satisfying the same properties, arguing that Shannon entropy can be taken as a robust parameter-free default.
+3.2 Entropy changes as a proxy for expansion and pruning strategies
+By definition, entropy is maximized when the probability mass is equally distributed, and minimized when it is concentrated in one event.
+In our setting, we focus on the entropy of layer-wise next-token predictions extracted via logit-lens (Section 4).
+Beyond its absolute value, we also analyze differences in the entropy of intermediate predictions between a layer $i$ and its predecessor $\Delta H_{i}=H_{i}-H_{i-1}$.
+We argue that we can interpret these increments and reductions as strategies of expansions and pruning of the candidate set, respectively.
+More formally, we define a candidate to be a token appearing in the top-$p$ distribution, with $p=0.6$.
+To this end, we need to validate two claims:
+  1. C1
+    $\Delta H_{i}$ is monotonically related to changes in the number of possible candidates: positive (negative) entropy variations correspond to increases (decreases) in the number of candidates, with larger variations indicating larger changes.
+  2. C2
+    Distributions across layers share a non-negligible portion of their candidates.
+Claim C1 ensures that variations in entropy, $\Delta H$, reflect changes in the size of the top-$p$ candidate set, rather than fluctuations confined to the bottom-(1p)(1-p) tail of the distribution.
+Claim C2 ensures that the top-$p$ candidate sets remain largely stable across layers.
+Together, these two claims allow us to interpret increases and decreases of entropy as expansion and pruning of the next-token candidate set.
+We validate C1 by measuring the correlation between $\Delta H_{i}$ and the change in the number of candidates between layer $i$ and its predecessor.
+Instead of using Pearson correlation, which assumes a linear relationship, we compute the Spearman rank correlation (Gibbons, 1993).
+This allows us to test for a monotonic relationship between entropy variations and changes in the number of top-$p$ candidates, without assuming linearity.
+Table 1 shows a high correlation, confirming the validity of C1.
+We give further details in Appendix A.1.
+Table 1: Spearman rank correlation between layer-wise entropy variations ($\Delta H_{i}$) and changes in the number of next-token candidates across layers, confirming that $\Delta H$ captures meaningful candidate expansion and pruning (Claim C1).
+Model Spearman Correlation
+Llama-3.2-1B 0.8250
+Llama-3.2-3B 0.7365
+Gemma-2-2B 0.8804
+Gemma-2-9B 0.8540
+To validate C2, we measure the percentage of top-$p$ tokens shared by layer $i$ and its predecessor. Figure 1 shows that distributions across layers share many of their candidates, confirming C2.
+In summary, the validation of these claims grounds our analysis in a concrete mechanism: it allows us to interpret increments and reductions of entropy as strategies of expansion and pruning of the candidate set, respectively.
+This justifies the use of $\Delta H$ as an interpretable, low-dimensional view of the token prediction dynamics across layers and for individual generations.
+Figure 1: Fraction of top-$p$ ($p=0.6$) next-token candidates shared between a layer and its successor across different models.
+The consistently high overlap confirms that candidate sets remain stable across layers (Claim C2).
 3.3 Three regimes of Rényi entropy
+In this subsection, we show that our analysis is not tied to an arbitrary entropy definition, but extends to a broad class of dimensionality reductions for categorical distributions.
+Through the main body of the paper, we instantiate our framework using Shannon entropy.
+More generally, our analysis applies to the broader family of Rényi entropies $H_{\alpha}$, described in Section 2, which subsumes Shannon entropy as the special case $\alpha\rightarrow 1$.
+Empirically, we identify an informative regime of $\alpha$ values within which entropy profiles are stable and expressive; importantly, this regime includes the Shannon case $\alpha\rightarrow 1$.
+In Appendix A.5 and Section 5.3 (Table 3), we show that qualitative and quantitative results remain equivalent within this regime.
+We therefore adopt Shannon entropy as a robust default for the remainder of the paper, avoiding the need to tune additional hyperparameters.
+4 Entropy-Lens
+In LLMs, each block operates on the residual stream to map input token sequences to output token distributions.
+To study their evolution, we introduce Entropy-Lens, a simple, scalable, and model-agnostic framework that summarizes layer-wise predictions through their entropy.
+As shown in Section 3.2, entropy provides a compact proxy for the number of next-token candidates at each stage of computation.
+Conceptually, Entropy-Lens can be seen as a dimensionality reduction of transformer computations, compressing complex layer-by-layer predictions into an entropy profile that provides a compact representation of how the model progressively expands and prunes its space of plausible continuations.
+In this section, we detail our methodology (further details in Appendix D.1).
+Entropy profiles.
+Formally, let $S=(t_{1},\dots,t_{N})$ denote an input sequence of tokens $t_{j}$ and let $x_{j}^{i}$ be the residual-stream activation of token $t_{j}$ after layer $i$.
+Similarly to nostalgebraist (2020), decoding this activation through the model’s output head and applying a softmax yields an intermediate probability distribution
+$y_{j}^{i}=W(x_{j}^{i}),\quad W:=\operatorname{softmax}\circ D$ (3)
+where $D$ is the decoder matrix. We then compute the entropy of this distribution $H(y_{j}^{i})$ to produce a single scalar per layer and token $h_{j}=(H(y_{j}^{1}),\dots,H(y_{j}^{L}))$.
+Repeating this computation across layers yields a sequence of entropy values which we call entropy profile.
+These profiles compress high-dimensional and unordered token prediction dynamics, providing a low-dimensional, layer-wise information-theoretic representation of the model’s computation.
+Crucially, this construction does not rely on gradients, fine-tuning, or auxiliary probes (Tang et al., 2024; Zhao et al., 2025; Zhang et al., 2026), and can be applied to frozen, off-the-shelf models, in a model agnostic way.
+Aggregated entropy profiles.
+While the entropy profile tracks the candidate pool for a single token, individual predictions are highly variable (Holtzman et al., 2020; Lin et al., 2024).
+Therefore, model behavior is also analyzed at the level of prompts or generations consisting of multiple tokens.
+To obtain a stable and informative representation at this level, entropy profiles must be aggregated across tokens.
+Specifically, given a prompt, we generate $T$ consecutive tokens autoregressively and extract the profile for each step, yielding a collection of entropy profiles $\{h_{1},\dots,h_{T}\}$.
+These profiles can be combined using simple aggregation operators to form a single representation of the model’s computation for the entire generation.
+In our experiments, we primarily use concatenation across tokens, resulting in a matrix-valued representation that preserves the layer-wise structure of entropy evolution for each generated token.
+Alternative aggregation strategies, such as averaging across tokens, are also supported.
+However, we focus on concatenation as it retains maximal temporal information while remaining straightforward to interpret.
+Figure 2: Overview of the Entropy-Lens pipeline.
+Left: intermediate residual-stream activations at each layer are projected into the output space via logit-lens, yielding layer-wise next-token distributions whose entropy is computed.
+Right: the resulting per-token entropy profiles are aggregated across generated tokens (e.g., by concatenation or averaging) to produce a representation that captures how the model expands and prunes its space of plausible continuations across depth and can optionally be fed to a classifier to assess profile distinctiveness.
+In the following sections, we use Entropy-Lens to show that entropy profiles make visible systematic patterns in how transformer models expand and contract their candidate pool across depth.
+These token prediction dynamics exhibit consistent structure within model families (Section 5.1), and vary across tasks (Section 5.2) and output formats (Section 5.3).
+Furthermore, the expansion and pruning phases prove to have different relative importance for downstream task performance (Section 5.4).
+5 Exploring prediction dynamics in LLMs
+The entropy profiles introduced in Section 4 provide a compact view of how intermediate next-token predictions evolve across depth, highlighting phases of expansion and pruning of the candidate set.
+In this section, we use Entropy-Lens to study the token prediction dynamics in LLMs.
+We show that different families, tasks and output formats are associated with distinct patterns of expansion and pruning of possible next-tokens.
+To analyze the differences in entropy profiles, we employ a k-nearest neighbors (kNN) classifier operating directly on the aggregated entropy representations defined in Section 4.
+The kNN is not used here as a proposed predictive model, but as a diagnostic tool: because a kNN relies entirely on distances between the inputs, its performance provides a direct indication of the distinctness of entropy profiles.
+Across all experiments in this section, entropy profiles are extracted from frozen models.
+Unless otherwise stated, we evaluate classification performance using standard cross-validation protocols and report area under the ROC curve (AUC) scores.
+Additional implementation details and experiments are provided in Appendix B and A.2, respectively.
+In the next subsections, we aim to answer the following research questions (RQs):
+  1. RQ1
+    Do different models combine expansion and pruning strategies differently in their token prediction dynamics? Do similarities in these dynamics correspond to structural similarities across models?
+  2. RQ2
+    Do these dynamics change depending on the task?
+  3. RQ3
+    Do these dynamics change depending on the output format?
+  4. RQ4
+    Given a particular model, what is the relative importance of one strategy over the other?
+5.1 Family specific dynamics (RQ1)
+We first use Entropy-Lens to analyze token prediction dynamics across models, assessing whether observed similarities align with structural similarities.
+Entropy profiles exhibit family-level patterns.
+We compare entropy profiles across 12 decoder-only language models (GPT, Gemma, Llama, Qwen) spanning a wide range of parameter counts (100M to 9B) and depths.
+Unless otherwise stated, we use a blank prompt and average over multiple generated tokens to obtain a stable profile for each model (details in Appendix B.1).
+Figure 3 reveals a striking pattern: entropy profiles cluster by model family rather than by size.
+In a low-dimensional projection, models from the same family form coherent groups, while distinct families remain well-separated.
+This indicates that prediction dynamics are not idiosyncratic to individual checkpoints, but reflect a shared computational structure induced by architectural and training choices.
+Qualitatively, different families exhibit distinct entropy regimes, as exemplified in Figure 4.
+GPT models start with high entropy and gradually sharpen their predictions, while Llamas display low-entropy initial stages followed by extended high-entropy plateaus and a final refinement.
+These differences suggest that model families combine expansion and pruning strategies in systematically different ways.
+Depth-rescaling invariance and coarse-to-fine behavior.
+A remarkable regularity emerges when comparing models of different scales in the same family.
+While absolute depth varies substantially, entropy profiles exhibit a striking alignment when mapped to a relative layer index (Figure 4).
+This suggests that model scale primarily dictates the resolution of the computation rather than its qualitative form.
+From this perspective, smaller models (less than 4B parameters) appear to implement a coarser discretization of the token prediction dynamics found in their larger counterparts: additional layers serve to refine the signal without fundamentally altering the underlying sequence of expansion and pruning.
+Overall, these findings establish entropy profiles as readable fingerprints of model-specific token prediction dynamics.
+Figure 3: t-SNE projection of aggregated entropy profiles obtained from a blank prompt for different model families.
+Profiles cluster by family rather than by model size, indicating family-specific token prediction dynamics.
+Point size scales with number of parameters.
+Figure 4: Average entropy profiles over 32 generated tokens for different model families.
+When depth is normalized, models within the same family exhibit aligned entropy trajectories, revealing family-specific coarse-to-fine token prediction dynamics.
+5.2 Task specific dynamics (RQ2)
+We now use our framework to assess whether different task types have characteristic token prediction dynamics.
+We consider three task types that differ in the nature of computation they require: generative (continue a story), syntactic (count the number of words), and semantic (extract the main idea/subject).
+Dataset and prompts.
+We build prompts using the TinyStories dataset (Eldan and Li, 2023).
+For each task type, we construct prompts by combining a task instruction with a story, using three templates designed to reduce surface-level confounds: Base (task prompt + story), Reversed (story + task prompt), and Scrambled (either task prompt + scrambled story or scrambled story + task prompt), where a scrambled story is obtained by randomly permuting the words in a story.
+The scrambled condition acts as a control to verify that the entropy signature is driven by the task’s semantic and syntactic requirements rather than simple lexical patterns.
+Additionally, to further control for prompt phrasing, we use two semantically equivalent variants for each task instruction,
+as detailed in Appendix B.2.
+Protocol.
+We generate 800 prompts per task type (2400 total), balanced across templates.
+For each prompt, we extract aggregated entropy profiles as in Section 4 and evaluate task-type separability using a kNN classifier as a diagnostic probe.
+We report one-vs-rest ROC-AUC under 10-fold cross-validation.
+Results.
+Across all evaluated models, Table 2 shows that the kNN achieves high AUC, indicating that entropy profiles, and thus token prediction dynamics, present task-relevant structure.
+Appendix B.2 gives additional details and visualizations, while Appendix A.3 shows the benefits of tracking the entire entropy profile.
+Table 2: Task-type classification results on TinyStories using kNN over aggregated entropy profiles.
+High ROC-AUC scores indicate that entropy and token prediction dynamics exhibit task-specific characteristics.
+Model Size k-NN AUC (3 classes)
+Gemma-2-it 2.1B 97.66 $\pm$ 0.47
+Gemma-2-it 8.9B 98.38 $\pm$ 0.50
+Llama-3.2-it 1B 94.94 $\pm$ 0.79
+Llama-3.2-it 3B 94.77 $\pm$ 0.93
+Llama-3-it 8B 96.10 $\pm$ 0.67
+Phi-3 3.6B 97.07 $\pm$ 0.87
+5.3 Output format specific dynamics (RQ3)
+Similarly to Section 5.2, we now use Entropy-Lens to assess whether different formats of generated text have characteristic token prediction dynamics, independently of topical content.
+We construct a topic-format dataset by prompting models to generate short texts across multiple topics while enforcing one of three different formats among poem, scientific piece and chat log.
+Dataset construction.
+Each prompt is formed by pairing a format instruction with a topic, as described in Appendix B.4.
+For each generation, we extract aggregated entropy profiles over the generated text according to Section 4.
+Protocol and results.
+We evaluate format separability using the same kNN diagnostic setup as above and report ROC-AUC under 10-fold cross-validation.
+Entropy profiles exhibit format-specific characteristics, enabling reliable discrimination between output formats across models (Table 3).
+We further find that the results remain stable when replacing Shannon entropy with Rényi entropies across a broad range of $\alpha$ values.
+To visualize similarities, we project aggregated profiles via PCA.
+The clear clustering in Figure 5 demonstrates linear separability, confirming that format-specific computation leaves a distinctive signature in the token prediction dynamics.
+Table 3: Format classification results on the topic–format dataset using kNN over aggregated entropy profiles, computed with different Rényi $\alpha$ values (with Shannon entropy corresponding to $\alpha=1$).
+Performance is stable across $\alpha$, and high ROC-AUC scores indicate that different output formats are associated with characteristic entropy and token prediction dynamics.
+Model $\boldsymbol{\alpha}$ k-NN AUC (3 classes)
+0.5 97.3 $\pm$ 1.6
+Gemma-2-2B-it 1.0 98.7 $\pm$ 1.1
+5.0 98.4 $\pm$ 1.7
+0.5 97.8 $\pm$ 1.6
+Llama-3.2-1B-it 1.0 97.8 $\pm$ 2.4
+5.0 96.6 $\pm$ 2.6
+Overall, we have shown that Entropy-Lens proves to be a meaningful tool for the dimensionality reduction of the categorical distributions produced at each transformer layer.
+Moreover, it enables the analysis of how models expand and prune the set of candidate continuations, revealing shared token prediction dynamics within architectural families and distinct dynamics across tasks and output formats.
+In the next section, we study how forcing a model to alter its dynamics affects its performance.
+Figure 5: PCA projection of aggregated entropy profiles extracted from the topic–format dataset.
+Profiles cluster by output format, indicating that different formats are associated with characteristic entropy and token prediction dynamics.
+5.4 Intervening on token prediction dynamics (RQ4)
+Figure 6: Change in MMLU accuracy when skipping fractions of transformer blocks selected by maximal entropy increase (expansion, $\max\Delta H$), maximal entropy decrease (pruning, $\min\Delta H$), or at random.
+Disrupting expansion phases is generally more detrimental to downstream accuracy, with a notable exception in Gemma2-2B-it.
+Entropy-Lens naturally decomposes token prediction dynamics into two strategies: candidate expansion and pruning (as discussed in Section 4).
+In this section, we assess the relative importance for downstream accuracy of these two strategies by selectively disrupting layers associated with maximal entropy increases ($\max\Delta H$) and decreases ($\min\Delta H$), respectively.
+We conduct this analysis in a multiple-choice question setting using the MMLU benchmark (Hendrycks et al., 2021).
+The choice of this dataset is due to several convenient qualities: its strong fact-based components retain good performance on base models, and it possesses a mild reasoning component, essential in a 1-shot setting.
+We consider six LLMs, including both base and instruction-tuned (it) variants: Gemma2-2B-base, Gemma2-9B-base, Gemma2-2B-it, Gemma-9B-it, Llama3.2-3B-base and Llama3.2-1B-it.
+Setup.
+For each MMLU question, we evaluate the LLM’s prediction by the highest probability token among {A, B, C, D} and comparing it to ground truth.
+To probe the functional roles of expansion and pruning, we identify the top-$k$ layers that maximize or minimize $\Delta H$ per question.
+These layers are skipped during inference to measure the impact on accuracy.
+We vary $k$ from 10% to 50% of the total layers in 10% increments.
+We employ random layer skipping as a baseline to control for the performance degradation caused by model truncation.
+This allows us to determine if our targeted strategies yield impacts beyond simple architectural degradation.
+Furthermore, a random baseline avoids the biases inherent in sequential or fixed-order deletion strategies.
+In this setting, skipping layer $i$ is implemented by setting the contribution of the $i$-th transformer block to the residual stream to zero.
+Results.
+Figure 6 shows that expansion ($\max\Delta H$) and pruning strategies ($\min\Delta H$) have different relative impact for downstream performance.
+Across all Llama variants, we observe a critical dependency: skipping the single layer corresponding to maximal expansion causes performance to collapse to random chance.
+For Gemma2-2B-Base, Gemma2-9B-Base, and Gemma2-9B-it, we observe a distinct separation between the effects of disrupting expansion versus pruning phases, with the former being more detrimental.
+In contrast, Gemma2-2B-it exhibits a notable inversion of this dynamic, where pruning phases prove significantly more critical for downstream accuracy than expansion phases.
+Whether this reversal in functional relevance is a direct byproduct of specific alignment or fine-tuning mechanisms remains an open question for future work.
+6 Related Work
+Lenses in LLMs.
+Mechanistic interpretability (Bereska and Gavves, 2024) aims to provide a precise description and prediction of transformer-based computations.
+Common tools in the field are lenses, which are a broad class of probes deployed in intermediate steps of the residual stream.
+For example, logit-lens (nostalgebraist, 2020) uses the model’s decoder to project the intermediate activations in the vocabulary space.
+Tuned-lens (Belrose et al., 2023) refines this technique by training a different affine probe at each layer, instead of only using the pretrained model’s decoder.
+Our framework, Entropy-Lens, builds on the Transformer-Lens library (Nanda and Bloom, 2022), which employs logit-lens to study and characterize LLMs’ computations via their decoded version with information theory.
+Information theory in LLMs.
+Information theory has been applied to both LLM training and interpretability.
+For instance, attention entropy collapse is linked to training instabilities (Zhai et al., 2023), and matrix entropy evaluates ‘compression’ (Wei et al., 2024; Queipo-de-Llano et al., 2025).
+In a similar vein, Skean et al.
+(2024, 2025) adapt metrics such as prompt entropy and curvature to evaluate the quality of intermediate representations.
+Additionally, mutual information helps quantify chain-of-thought effectiveness (Ton et al., 2024).
+In contrast, Entropy-Lens shifts focus to the vocabulary domain, analyzing the entropy evolution of intermediate decoded logits.
+While Dombrowski and Corlouer (2024) use similar measures to distinguish truthful from deceptive generations under explicit instruction, we uncover broader signatures across model families, tasks, and formats.
+Regarding memorization, Brown et al.
+(2021) used Shannon mutual information between data and models, while Morris et al.
+(2025) extended this using instance-level Kolmogorov theory (see Appendix C.1 for a derivation linking our entropy measure to memorization).
+7 Conclusion
+Most interpretability research focuses on internal latent representations, often leaving the high-dimensional and categorical dynamics of the token space underexplored.
+To bridge this gap, we introduced Entropy-Lens, a framework that distills the token prediction dynamics of the residual stream in token space into a low-dimensional signal: the entropy profile.
+By applying this metric, we showed that the entropy of logit-lens predictions overcomes the challenges of standard statistical descriptors, providing a per-layer, permutation-invariant scalar that uncovers token prediction dynamics driven by distinct expansion and pruning strategies (Section 3).
+Our empirical evaluation across various model sizes and families demonstrates that token prediction dynamics are structured computational signatures.
+We found that they are family-specific and invariant under depth rescaling (Section 5.1), suggesting that larger models within a family refine the same underlying decision process.
+Furthermore, we observed that expansion and pruning strategies are combined in task- and format-dependent ways, giving rise to token prediction dynamics that are characteristic of task type (Section 5.2) and output format (Section 5.3).
+Finally, our intervention experiments revealed that these strategies have unequal impact on downstream performance, with the expansion strategy usually being more critical for maintaining accuracy (Section 5.4).
+Ultimately, these findings enhance our understanding of the residual stream, enabling a granular assessment of information processing across model depth.
+Limitations and future work.
+Despite the insights provided by Entropy-Lens, our study has several limitations that outline directions for future work.
+First, while we established that entropy profiles are characteristic of model families, we lack a mechanistic understanding of which specific architectural components or training schemes play the causal role in shaping a particular entropy profile.
+Second, regarding the functional role of these strategies, it remains unclear whether and how different training stages, such as fine-tuning or RLHF, alter the relative importance of expansion and pruning phases for downstream performance.
+Finally, our analysis has been confined to decoder-only Transformers; we have not yet extended this framework to architectures with different information flow paradigms, such as ViTs (Appendix A.4) or Mixture-of-Experts.
+Acknowledgements
+We thank Lorenzo Sani for insightful discussions and for his help in reviewing the paper prior to submission.
+Impact statement
+This paper presents work whose goal is to advance the field of machine learning.
+There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
+Appendix A Additional Experiments
+To further test the generality and flexibility of our methodology, we conduct additional experiments beyond the core settings presented in the main text.
+In particular, we explore how Entropy-Lens performs in a different modality: computer vision.
+Without any architectural adjustment or fine-tuning, we apply the same framework to Vision Transformers (ViTs) and observe that entropy profiles extracted from visual models exhibit qualitatively interpretable structure.
+These preliminary results suggest that our method may extend beyond language models, but a systematic evaluation across modalities is left for future work.
+We further performed a complementary study on the TinyStories dataset, using the Gemma-2-2b-it model, to assess which transformer blocks are most critical for task classification and whether all layers are necessary.
+In this setting, we compared full entropy profiles with reduced variants obtained from single layers or from equidistant subsets of layers (first, middle, and last).
+Our results show that the complete entropy profile achieves substantially higher classification accuracy, indicating that information is distributed across depth and cannot be captured by a small subset of layers alone.
+A.1 Correlation between changes in entropy and candidate counts
+Figure 7 demonstrates the relationship between changes in entropy and candidate counts in an example.
+The reported metrics are calculated by prompting the model with: "Q: Was Claude Shannon born in an even or odd year?
+A:"
+Figure 7: Relationship between changes in entropy and changes in candidate counts.
+Each box corresponds to a layer, colored in green if the candidate count increases, and red if it decreases.
+Inside each box, we report the value for $\Delta H$ of the corresponding layer.
+A.2 Entropy profiles contain signal about answer correctness
+We finally test whether entropy profiles contain information associated with answer correctness.
+We use the Massive Multitask Language Understanding (MMLU) benchmark (Hendrycks et al., 2021), which consists of multiple-choice questions spanning 57 subjects and varying difficulty levels.
+Setup.
+We evaluate two instruction-tuned models (one from the Llama family and one from the Gemma family) on MMLU.
+To control for prompt framing, we present each question using three prompting styles: Base, Instruct, and Humble (Appendix B.3).
+For each instance, we extract aggregated entropy profiles from the model generation and label them as correct/incorrect based on whether the selected option matches the ground truth.
+Protocol.
+We apply the kNN diagnostic classifier to distinguish correct vs.
+incorrect profiles and report ROC-AUC under 10-fold cross-validation.
+To account for class imbalance, we also compare performance against a dummy baseline that samples predictions according to the empirical class proportions.
+Results.
+Across prompting styles and both models, entropy profiles yield ROC-AUC values substantially above the baseline (Table 4), indicating that correctness-related signal is present in the entropy dynamics.
+Table 4: MMLU correctness for different models. The table reports both the LLM overall accuracy on MMLU and the kNN accuracy in distinguishing correct and wrong answers
+Model Prompt LLM-Acc. kNN AUC
+Base 50.89 73.61 $\pm$ 1.52
+Llama Humble 58.51 69.90 $\pm$ 1.06
+Instruct 60.62 67.23 $\pm$ 1.62
+Base 56.10 71.88 $\pm$ 1.63
+Gemma Humble 54.71 72.78 $\pm$ 1.15
+Instruct 56.38 68.36 $\pm$ 1.23
+A.3 Entropy profiles from different blocks
+To assess whether entropy profiles from all transformer layers are necessary for model characterization, or if comparable results can be achieved using fewer layers, we conducted an evaluation using different layer subsets.
+Specifically, we repeated the TinyStories experiments (Section 5.2) using four different configurations: (1) first layer only, (2) middle layer only, (3) last layer only, and (4) a combination of first, middle, and last layers.
+We then compared the classification accuracy of k-NN classifiers trained on these reduced entropy profiles against those using complete layer sequences.
+The results are visible in Table 5.
+Experimental setup.
+The k-NN classifier was configured with $k=11$ neighbors using Euclidean distance as the similarity metric.
+For sequence generation, we employed a sampling-based approach rather than deterministic decoding which was used for the results reported in the Section 5.2.
+Table 5: kNN AUC across different sets of considered layers for Gemma-2-2B-it. The results show the benefits of considering the entire entropy profile
+Considered layers kNN AUC
+[…]
+[…]
+[…]
+[…]
+[…]
+A.4 Entropy-Lens for Vision Transformers
+To demonstrate the versatility and robustness of our approach beyond language modeling, we give preliminary results on ViTs and DeiTs.
+Using 20 classes from ImageNet-1K (Russakovsky et al., 2015), with 20 images per class, and without any modifications to our framework, we generate the entropy profiles shown in Figure 8.
+We observe that all profiles start with high entropy values, which then decrease, mostly in the final layers.
+This behavior is qualitatively similar to that of GPTs or larger LLaMa models (Section 5.1), pointing to a possible common trend across domains as different as image processing and natural language processing.
+Focusing on computer vision models, we note that while ViT and DeiT families exhibit qualitatively similar trends, they differ quantitatively—ViTs start with higher entropy values, making them easily distinguishable from DeiTs.
+Notably, the only profile that stands out is that of ViT Large (with $\sim 300$M parameters), compared to the other models analyzed in this section, which have $\leq 86$M parameters.
+For ViT Large, entropy decreases more smoothly, appearing not only as a better approximation of the sharp drop seen in smaller models but possibly following a different behavior entirely, with the entropy decline starting earlier.
+We hypothesize a phase transition in entropy behavior as model size increases, occurring somewhere between 8787M and 307307M parameters, though a more extensive study would be required to confirm this hypothesis.
+Figure 8: Entropy profiles for ViT model families.
+A.5 Stability across Rényi entropy variants
+To qualitatively explore how the Rényi entropy affects the structure of entropy profiles, we compute pairwise cosine similarities between profiles generated with different values of $\alpha$.
+This analysis is performed on a subset of the topic-format dataset (see Appendix B.4), and the resulting similarity matrices are shown in Figure 9.
+We observe three distinct regimes as $\alpha$ varies: (1) For very small values of $\alpha$ (e.g., $\alpha<0.2$), the similarity matrices are nearly flat, with profiles being almost identical across examples.
+This is expected, as Rényi entropy in this regime weights all tokens with non-zero probability almost equally, and usually all tokens have non-null probability.
+(2) For large values of $\alpha$ (e.g., $\alpha>20$), the similarity matrices also flatten.
+In this case, the entropy becomes increasingly dominated by the few tokens with highest probabilities.
+Since these sets of top tokens tend to have similar cardinalities (in the limit equal to 1), the profiles collapse into a narrow set of values, losing expressiveness and becoming more sensitive to local fluctuations.
+(3) Between these extrema lies an informative regime, approximately $0.5\leq\alpha\leq 20$, where entropy profiles are heterogeneous enough to retain meaningful variation.
+This is reflected in the standard deviation of the similarity matrices, which peaks in this interval.
+We find that the qualitative structure of entropy profiles, shown in Figure 9, and the diagnostic results reported in Section B.4 are stable when replacing Shannon entropy with Rényi entropies for a broad range of $\alpha$ values.
 Figure 9: Cosine similarity matrices between entropy profiles computed on a subset of the topic-format dataset using different values of $\alpha$ in the Rényi entropy.
@@
 Note how similarity flattens for very low and very high $\alpha$, while intermediate values yield more informative profiles.
-To qualitatively explore how the Rényi entropy affects the structure of entropy profiles, we compute pairwise cosine similarities between profiles generated with different values of $\alpha$.
-This analysis is performed on a subset of the topic-format dataset, and the resulting similarity matrices are shown in Figure 7.
-We observe three distinct regimes as $\alpha$ varies:
-  • For very small values of $\alpha$ (e.g., $\alpha<0.2$), the similarity matrices are nearly flat, with profiles being almost identical across examples.
-This is expected, as Rényi entropy in this regime weights all tokens with non-zero probability almost equally, and usually all tokens have non-null probability.
-  • For large values of $\alpha$ (e.g., $\alpha>20$), the similarity matrices also flatten.
-In this case, the entropy becomes increasingly dominated by the few tokens with highest probabilities.
-Since these sets of top tokens tend to have similar cardinalities (in the limit equal to 1), the profiles collapse into a narrow set of values, losing expressiveness and becoming more sensitive to local fluctuations.
-  • Between these extremes lies an informative regime—approximately $0.5\leq\alpha\leq 20$—where entropy profiles are heterogeneous enough to retain meaningful variation.
-This is reflected in the standard deviation of the similarity matrices, which peaks in this interval.
-This qualitative observation supports our empirical findings: in Section 5.5, we show that format classification accuracy remains high and stable within this informative $\alpha$ range.
-Notably, Shannon entropy ($\alpha=1$) falls within this interval, providing a strong justification for its use in the main experiments.
-By choosing $\alpha=1$, we retain discriminative power while avoiding the need to tune additional hyperparameters.
-5.5 Entropy profiles identify text format
-We test whether the entropy profiles contain signal about the format of the generated text.
-In this experiment, we prompt the model to produce short texts across different topics while enforcing one of three predefined formats: poem, scientific piece, or chat log.
-We call these generated texts the topic-format dataset.
-Model $\alpha$ k-NN ROC-AUC
-Gemma-2-2b-it 0.5 97.3 ± 1.6
-1.0 98.7 ± 1.1
-5.0 98.4 ± 1.7
-Llama-3.2-1b-it 0.5 97.8 ± 1.6
-1.0 97.8 ± 2.4
-5.0 96.6 ± 2.6
-Table 2: ROC-AUC (one-vs-rest) results of two models with three values of $\alpha$ for the Rényi entropy for format classification.
-Standard deviations are computed over 10-fold cross-validation splits.
-We use entropy profiles computed not only with Shannon entropy ($\alpha=1$), but also with Rényi entropy for $\alpha\in\{0.5,5.0\}$.
-As shown in Table 2, the k-NN classifier achieves high ROC-AUC scores across all values of $\alpha$, indicating that the format is reliably encoded in the entropy profiles.
-Moreover, the small variability in performance across $\alpha$ values supports our use of Shannon entropy as a principled, parameter-free default.
-To better understand the structure of the entropy profiles, we perform Principal Component Analysis (PCA) and visualize the first two components.
-As shown in Figure 6, the profiles cluster distinctly by format, forming linearly separable groups in the reduced space.
-This indicates that format-specific computation patterns are not only detectable by a classifier, but visibly reflected in the global shape of the entropy evolution across layers.
-Model Prompt LLM Acc. k-NN AUC-ROC
-Llama Base 50.89 73.61 ± 1.52
-Humble 58.51 69.90 ± 1.06
-Instruct 60.62 67.23 ± 1.62
-Gemma Base 56.10 71.88 ± 1.63
-Humble 54.71 72.78 ± 1.15
-Instruct 56.38 68.36 ± 1.23
-Table 3: Performance comparison of two models (Llama-3.2-3B-it and Gemma-2-2B-it) for different prompt types on MMLU accuracy and k-NN AUC-ROC score.
-6 Conclusions
-In this work, we prototyped a novel model-agnostic interpretability framework for large-scale transformer-based architectures grounded in information-theory.
-In Section 5.1, we showed that the entropy profiles of LLM intermediate predictions identify the LLM’s model family.
-Additionally, in Sections 5.2 and 5.5 we showed that the same entropy profiles can be used to identify the ‘task type’ and the ‘format type’ in LLMs, whilst in Section 5.3 we used them to distinguish between correct and wrong LLM generated answers.
-Moreover, we investigated the validity of using Shannon entropy instead of the more general Rényi entropy, and empirically demonstrated its effectiveness.
-Importantly, all our experiments were conducted on frozen off-the-shelf large-scale transformers.
-6.1 Limitations and Future Work
-While this work paves the way to further investigations in information theoretic interpretability, it also presents a number of limitations.
-First, the concepts of ‘task type’ and ‘format type’ don’t have a formal and well established definition.
-Moreover, we showed how different models possess different characteristic entropy profiles.
-We conjecture that these particular shapes are a byproduct of training procedure and architectural designs, but future research could focus on understanding the precise connections.
-Currently, we do not have a theoretical explanation for the observed phenomena or their underlying causes.
-Another interesting line of research could focus on considering more fine-grained measures of information instead of just an aggregated one such as entropy.
-With these limitations in mind, our methodology could be used to probe the reasoning capabilities of LLMs, for instance by comparing the entropy profile of a reasoning task vs.
-a data retrieval task.
-If these profiles happen to match, it could be taken as an argument against the ability of LLMs to reason.
-Conversely, if they do not, it may suggest that some form of reasoning is indeed occurring.
-Finally, recent literature explored the use of entropy for private inference (PI), where computations are performed on encrypted data without revealing users’ sensitive information (Jha and Reagen 2025).
-While previous work focused on the entropy of the attention mechanism, future research could use our methodology to develop PI-friendly applications of LLMs.
-Appendix A Appendix Structure
-The appendix is organized as follows:
-  • Appendix B – Additional Experiments: We assess the generality of our approach by applying it to Vision Transformers (ViTs).
-This analysis demonstrates that entropy profiles can be extracted and interpreted even in non-language domains, without any modification to our framework.
-  • Appendix C – Evaluation Details: We provide full details on the datasets and prompt templates used in our experiments.
-In particular, we specify how prompts were constructed for task classification, output correctness, and format classification tasks.
-This section also includes hardware information to support reproducibility.
-  • Appendix D – Minimal Implementation: We present a minimal code snippet that reproduces the core entropy profile extraction logic in a few lines of code.
-While our full codebase offers several optimizations and utilities, this section emphasizes transparency and ease of replication by showcasing the conceptual simplicity of our approach.
-Appendix B Additional Experiments
-To further test the generality and flexibility of our methodology, we conduct additional experiments beyond the core settings presented in the main text.
-In particular, we explore how Entropy-Lens performs in a different modality: computer vision.
-Without any architectural adjustment or fine-tuning, we apply the same framework to Vision Transformers (ViTs) and demonstrate that entropy profiles extracted from visual models exhibit meaningful and interpretable structure.
-This supports the claim that our method is not only scalable and model-agnostic, but also transferable across modalities.
-B.1 Entropy-Lens for Vision Transformers
-To demonstrate the versatility and robustness of our approach beyond language modeling, we analyze the entropy profiles of ViTs and DeiTs.
-Using 20 classes from ImageNet-1K (Russakovsky et al.
-2015), with 20 images per class, and without any modifications to our framework, we generate the entropy profiles shown in Figure 8.
-We observe that all profiles start with high entropy values, which then decrease, mostly in the final layers.
-This behavior is qualitatively similar to that of GPTs or larger LLaMa models (Section 5.1), suggesting a universal pattern across domains as different as image processing and natural language processing.
-Focusing on computer vision models, we note that while ViT and DeiT families exhibit qualitatively similar trends, they differ quantitatively—ViTs start with higher entropy values, making them easily distinguishable from DeiTs.
-Notably, the only profile that stands out is that of ViT Large (with $\sim 300$M parameters), compared to the other models analyzed in this section, which have $\leq 86$M parameters.
-For ViT Large, entropy decreases more smoothly, appearing not only as a better approximation of the sharp drop seen in smaller models but possibly following a different behavior entirely, with the entropy decline starting earlier.
-We hypothesize a phase transition in entropy behavior as model size increases, occurring somewhere between 8787M and 307307M parameters.
-Figure 8: Entropy profiles for ViT model families.
-Appendix C Evaluation Details
+Appendix B Evaluation Details
 This section provides further details on the datasets and prompt templates used to evaluate the effectiveness of entropy profiles in the main experiments.
@@
 To this end, we employed multiple prompt variations.
-The subsections below report the full set of templates used, as referenced in Sections 5.2, 5.3, and 5.5 of the main paper.
+The subsections below report the full set of templates used, as referenced in Sections 5.2, A.2, and 5.3 of the main paper.
 The scripts used to generate these datasets—along with the full codebase to extract entropy profiles and reproduce all experiments—are shared as part of our code release.
 Finally, we also provide information about the hardware used to run our experiments to facilitate reproducibility.
-C.1 Prompt templates for TinyStories tasks
+B.1 Model family classification
+In Section 5.1, we show how entropy profiles can effectively identify both model families and model sizes.
+Our analysis reveals that entropy profiles exhibit qualitatively distinct patterns across different model families and sizes, as illustrated through scatterplot visualizations.
+By applying t-SNE dimensionality reduction, we cluster models by family, indicating that entropy profiles capture meaningful structural differences between architectures.
+To quantitatively assess the classification capabilities of entropy profiles, we employ a k-nearest neighbors classifier ($k=3$ and euclidean distance) to predict both model families and sizes based on their entropy traces.
+The classification results are presented in Table 6.
+To obtain labeled model size categories, we binned models into 4 classes based on parameter count (in billions): <<1B, 1-3B, 3-5B, and >>5B.
+Table 6: F1-scores for model family and model size classification. Each reported value is the mean across 10 runs, with the standard deviation computed over random 50/50 train–test splits.
+Task Macro F1-score
+[…]
+[…]
+Preprocessing Steps.
+Since entropy traces vary in length across models due to different layer counts, we apply linear interpolation to standardize all traces to the same length.
+Additionally, to ensure fair classification performance, we standardize the samples to reduce bias from scaling effects in the entropy profiles, allowing the classifier to focus on the characteristics of each trace.
+B.2 Prompt templates for TinyStories tasks
 In Section 5.2, we describe an experimental setup designed to test whether entropy profiles can identify different types of tasks.
-To this end, we used the TinyStories dataset (Eldan and Li 2023) and constructed prompts combining short stories with specific task instructions.
-Each task type—generative, syntactic, and semantic—was associated with two distinct natural language formulations, referred to as task prompts.
+To this end, we use the TinyStories dataset (Eldan and Li, 2023) and construct prompts combining short stories with specific task instructions.
+Each task type, generative, syntactic, and semantic, was associated with two distinct natural language formulations, referred to as task prompts.
 These are listed in Table 7.
 By varying the task prompt, we ensure that our classification results are not simply driven by surface-level textual artifacts, but instead reflect deeper computational signatures captured by the entropy profiles.
-This table complements the prompt templates (base, reversed, and scrambled) described in the main text and provides the full set of instructions used to elicit different model behaviors.
+The same table complements the prompt templates (base, reversed, and scrambled) described in the main text and provides the full set of instructions used to elicit different model behaviors.
+Figure 10 shows the mean and standard deviation of the entropy profiles for different tasks.
+Figure 10: Average entropy profiles with shaded standard deviation for different task types: generative, syntactic, and semantic.
+These tasks are induced with the prompts described in Appendix B.2.
+Left: Llama-3.2-it.
+Right: Gemma-2-it.
+Table 7: Prompt templates used for TinyStories tasks.
 Task Type Task prompt
@@
 Who is the subject of the story?
-Table 4: Prompt templates used for TinyStories tasks.
-C.2 Prompt templates for MMLU correctness classification
-In Section 5.3, we test whether entropy profiles can distinguish between correct and incorrect answers produced by language models.
-To construct the dataset for this experiment, we used the Massive Multitask Language Understanding (MMLU) benchmark (Hendrycks et al.
-2021), applying three distinct prompt styles to elicit different answer behaviors from the models.
+B.3 Prompt templates for MMLU correctness classification
+In Section A.2, we test whether entropy profiles can distinguish between correct and incorrect answers produced by language models.
+To construct the dataset for this experiment, we used the Massive Multitask Language Understanding (MMLU) benchmark (Hendrycks et al., 2021), applying three distinct prompt styles to elicit different answer behaviors from the models.
 Table 8 reports the full prompt templates used in this experiment.
@@
 This variation in prompting allows us to control for instruction framing and to evaluate whether entropy profiles can capture response confidence and correctness robustly across different model behaviors.
-The table shown here complements the description in the main text.
+Table 8 complements the description in the main text.
+Table 8: Prompt templates used for the MMLU dataset.
 Prompt Type Prompt
@@
 Answer:
-Table 5: Prompt templates used for the MMLU dataset.
-C.3 Prompt Construction for the Topic-Format Dataset
+B.4 Prompt Construction for the Topic-Format Dataset
 To evaluate whether entropy profiles captures stylistic features of generated text, we constructed a custom dataset referred to as the topic-format dataset.
@@
 This setup enables robust testing of the extent to which entropy profiles encode formatting cues, beyond topical content or task semantics.
-C.4 Experimental and hardware setup
-All experiments were conducted on a compute node equipped with an NVIDIA L40 GPU, an Intel Xeon Gold CPU, 128 GB of RAM, and running Ubuntu 22.04.
+B.5 Experimental and hardware setup
+The experiments were conducted on:
+  • A compute node equipped with an NVIDIA L40 GPU, an Intel Xeon Gold CPU, 128 GB of RAM, and running Ubuntu 22.04.
+  • A compute node equipped with an NVIDIA H100 GPU, an Intel Xeon Platinum CPU, 240 GB of RAM and running Ubuntu 22.04.
 The primary software frameworks used were PyTorch, Transformer-Lens, and HuggingFace Transformers.
 Inference on LLMs was performed using float16 precision for improved efficiency.
+Part of the experiments were conducted on Chameleon Cloud (Keahey et al., 2020).
+Source code is available at: github.com/christopher-irw/Entropy-Lens.
+Appendix C Theoretical Considerations
+In addition to the empirical results presented in the main text, we provide here a preliminary theoretical perspective that connects entropy profiles to the literature on memorization in transformers.
+Our goal is not to give a full formal treatment, but rather to outline how previous definitions of memorization, based on information theory, can be related to the quantities we compute.
+We first recall the frameworks introduced by Brown et al.
+(2021) and Morris et al.
+(2025), and then show how our entropy profiles can be interpreted as layer-wise estimates of memorization.
+C.1 Entropy and Memorization
+From Shannon to Kolmogorov memorization.
+Brown et al.
+(2021) introduced an information-theoretic framework to quantify memorization in trained models.
+Given a training data distribution $X$, a family of data-generating processes $\Theta$, and a training algorithm $L:X\mapsto\hat{\Theta}$ mapping training sets to trained models, they define memorization as the mutual information between $X$ and $\hat{\Theta}$
+[…]
+This quantity captures how much information about $X$ is retained in the distribution over trained models. It can be decomposed into
+$\text{mem}(X,\hat{\Theta})=\text{mem}_{I}(X,\hat{\Theta},\Theta)+\text{mem}_{U}(X,\hat{\Theta},\Theta),$ (5)
+where $\text{mem}_{I}$ measures generalization and $\text{mem}_{U}$ the unintended memorization (i.e., information about $X$ not attributable to the process $\Theta$).
+Building on this formalism, Morris et al.
+(2025) extended the analysis from distributions to individual instances, moving from Shannon to Kolmogorov information theory.
+The Kolmogorov complexity of an instance $x$ given model parameters $\hat{\theta}$ is
+$H^{k}(x|\hat{\theta})=\min_{s}\{|s|:f(s,\hat{\theta})=x\},$ (6)
+where $f$ is a computational model (imagine a decoder) that can take as input $x$ and $\theta$.
+The exact definition of Kolmogorov complexity is not computable in general.
+In practice, it can be approximated via arithmetic coding as
+[…]
+[…]
+Connecting to entropy profiles.
+[…]
+Now, rather than computing this value only for the full model, we do so for every intermediate truncation.
+Let $\hat{\Theta}_{i}$ denote the sub-model obtained by retaining only the first $i$ layers of the trained transformer and applying an early exit.
+The corresponding conditional entropies are
+$H(X|\hat{\Theta}_{i}),\quad i=1,\dots,N,$ (8)
+and the sequence $\{H(X|\hat{\Theta}_{i})\}_{i=1}^{N}$ constitutes the entropy profile. This is equivalent to
+$H(X|\hat{\Theta}_{i})=H(X)-I(X,\hat{\Theta}_{i}),$ (9)
+i.e. the negative mutual information between the dataset and the truncated model up to a constant $H(X)$, which is equal for all layers.
+This perspective suggests that entropy profiles can be interpreted as measuring how memorization is distributed across depth.
+Crucially, our empirical results show that this allocation does not follow a simple monotonic trend, as one might have expected a priori.
+Instead, it varies in a systematic way depending on model family, task, format, and confidence, revealing non-trivial patterns of information storage that had not been documented before.
 Appendix D Minimal Implementation
@@
 D.1 Minimal Entropy Profile Extraction
-The code in listing 1 demonstrates how to compute an entropy profile for a single prompt using a standard decoder-only transformer.
+The code in Listing LABEL:lst:entropy-signature demonstrates how to compute an entropy profile for a single prompt using a standard decoder-only transformer.
 It relies only on model forward passes and the use of logit-lens-style decoding.
-No gradients, fine-tuning, or access to internal model weights are required.
+No gradients or fine-tuning are required, but only forward access to intermediate hidden states and the output head.
 1from transformers import AutoTokenizer, AutoModelForCausalLM